### v1

# GT2-GS: Geometry-aware Texture Transfer for Gaussian Splatting

Abstract.
Transferring 2D textures to 3D modalities is of great significance for improving the efficiency of multimedia content creation. Existing approaches have rarely focused on transferring image textures onto 3D representations. 3D style transfer methods are capable of transferring abstract artistic styles to 3D scenes. However, these methods often overlook the geometric information of the scene, which makes it challenging to achieve high-quality 3D texture transfer results. In this paper, we present GT2-GS, a geometry-aware texture transfer framework for gaussian splitting. From the perspective of matching texture features with geometric information in rendered views, we identify the issue of insufficient texture features and propose a geometry-aware texture augmentation module to expand the texture feature set. Moreover, a geometry-consistent texture loss is proposed to optimize texture features into the scene representation. This loss function incorporates both camera pose and 3D geometric information of the scene, enabling controllable texture-oriented appearance editing. Finally, a geometry preservation strategy is introduced. By alternating between the texture transfer and geometry correction stages over multiple iterations, this strategy achieves a balance between learning texture features and preserving geometric integrity. Extensive experiments demonstrate the effectiveness and controllability of our method. Through geometric awareness, our approach achieves texture transfer results that better align with human visual perception. Our homepage is available at this https url.
Texture Transfer, 3D Gaussian Splatting, Geometry-aware
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
Figure 1. Previous methods struggle to accurately transfer complex texture appearances to the scene. In contrast, our GT2-GS framework injects geometric information into the scene during texture transfer, enabling geometry-aware texture transfer. Our method enables continuous and controllable texture transfer, and is capable of producing perspective effects as illustrated by the blue boxes.
1. Introduction
Texture transfer (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) aims to transfer the texture features of a reference image to a target image while preserving the structural and semantic content. Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis and were mainly applied to tasks such as image inpainting and data augmentation. With the advancement of deep learning, a variety of network-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Cao et al., 2023; Pu et al., 2024) have emerged. These methods enable texture transfer by incorporating semantic information, allowing for artistic content creation.
With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has grown significantly. Texture is an essential component of various graphics pipelines and plays a crucial role in producing high-quality 3D assets. Traditionally, texture creation requires skilled artists to design them manually, which is both costly and inefficient. Recently, the emergence of neural radiance fields (NeRF) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has greatly improved the efficiency of 3D content creation. Building upon such 3D representations, it is promising to develop 3D texture transfer method that directly transfers texture features from reference images to 3D representations. This approach bypasses the labor-intensive UV mapping process, significantly reducing both the cost and technical barriers of 3D content creation. Due to the challenge of multi-view consistency, 2D texture transfer methods (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) cannot be directly applied to 3D scenes. Some diffusion-based approaches (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) generate high-quality texture maps for 3D meshes. However, these methods modify the appearance of 3D objects via texture maps rather than directly transferring texture features into 3D representations. Moreover, they typically operate at the object level, making it difficult to generate textures for entire scenes. Therefore, how to achieve efficient 3D texture transfer remains an open and worthwhile research question.
Existing 3D scene stylization methods (Zhang et al., 2022; Liu et al., 2023, 2024; Zhang et al., 2024a; Liu et al., 2025) focus on transferring the artistic styles of abstract reference images to 3D scenes. We attempt to use texture images as references and apply state-of-the-art optimization-based stylization methods to transfer their texture features to 3D scenes. However, as shown in the first row on the left side of Fig. 1, the texture transfer results of this approach are suboptimal. Through analysis, we identify the core reason why existing 3D stylization methods cannot be directly applied to texture transfer. Texture is closely tied to geometry. 3D texture transfer needs to be harmonious with both the 3D geometry and the texture features of the reference image, requiring consistency between texture and geometry to ensure strict multi-view consistency. However, existing 3D stylization methods overlook the geometry of the scene. Firstly, 3D scenes are optimized via gradient descent using losses computed between rendered feature maps and texture features. Since texture images typically contain only a single type of texture, applying highly repetitive texture features across diverse rendered views is suboptimal. This fails to align with the rich geometric information in the rendered images, such as perspective geometry, edges, contours, and shapes. Moreover, current style transfer loss functions are computed independently across views, without considering the 3D geometric relationships between different viewpoints. This limits the ability to achieve spatially consistent and controllable texture transfer.
In this paper, we propose a novel framework GT2-GS for achieving geometry-aware texture transfer. Our insight is to integrate comprehensive geometric information into the texture transfer process to achieve controllable results that align with human visual perception. To enable geometry-aware texture transfer, we first propose a geometry-aware texture augmentation module. This module enhances texture features through scaling and rotation operations based on geometric information from the rendered image, such as perspective, edges, and contours. In addtion, a geometry-consistent texture Loss is proposed for 3D texture transfer. This loss uses the camera poses and depth geometry between a prior view and the current optimization view, achieving scene appearance editing with controllable texture orientation. To balance texture learning and geometry preservation, we further introduce a Geometry Preservation Strategy. This strategy alternates the optimization of Gaussian parameters between a texture transfer stage and a geometry correction stage, with multiple iterations performed. In this process, the geometry correction stage utilizes content images with ground truth for optimization, ensuring the accurate refinement of scene geometry. Extensive experiments demonstrate that our proposed framework achieves high-quality texture transfer results.
Our main innovations are as follows:
  • To the best of our knowledge, we are the first to incorporate geometric information into 3D texture transfer. The proposed framework enables direction-controllable and high-quality texture transfer results.
  • Geometry-aware texture Augmentation module and geometry-consistent texture loss are proposed to incorporate the geometric information of the scene during the 3d scene texture transfer. Moreover, we introduced a geometric preservation strategy to maintain the correct geometric information of the scene.
  • Extensive experiments demonstrate that, compared to SOTA methods, our proposed work achieves texture transfer results more consistent with human visual perception. Our proposed method maintains real-time rendering and multi-view consistency. The source code will be released.
2. Related Work
2.1. Texture Transfer
Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis. While preserving the structural information of the target image to achieve texture transfer effects. Gatys et al. (Gatys et al., 2016) pioneered the use of convolutional neural networks (CNNs) (Simonyan and Zisserman, 2014) for image style transfer, achieving impressive results. Subsequently, a variety of deep learning-based style transfer methods (Chen and Schmidt, 2016; Li and Wand, 2016; Huang and Belongie, 2017; Huo et al., 2021) have been proposed. These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer. Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (Vaswani et al., 2017), some deep learning-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Pu et al., 2024) can transfer local textures from the reference image to the target image based on high-level semantic correspondence. With the development of generative models (Goodfellow et al., 2014; Ho et al., 2020), some methods (Lyu et al., 2021; Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) have adopted texture images as guidance conditions for texture transfer. Notably, high-quality 3D texture maps can be directly generated using diffusion-based methods (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024). However, these methods require precise computation on existing meshes and are difficult to extend to complex and large-scale scenes. In this work, we explore how to directly transfer texture features into 3D representations.
Figure 2. Overview of GT2-GS. The overall pipeline is illustrated in (a). The input to our proposed framework includes the scene Gaussians, a content image, and a texture image. The texture image is transformed into a texture feature set through the GTA module, as shown in (b). This feature set is then used to construct the target feature map in the GT Loss, illustrated in (c). After embedding the parameter $c^{g}$, the Gaussians are optimized using the two-stage iterative strategy GPS, as shown in (d). As a result, the texture features are seamlessly integrated into the Gaussian representation.
2.2. 3D Style Transfer
3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency. Early research in 3D scene style transfer focused on explicit representations such as point clouds (Cao et al., 2020; Huang et al., 2021) ,meshes (Höllein et al., 2022; Yin et al., 2021), and voxels (Guo et al., 2021; Klehm et al., 2014). Zhang et al. (Zhang et al., 2022) explored the potential of NeRF in 3D stylization tasks. By optimizing the scene representation with the proposed 3d style transfer loss, it achieved pleasing visual effects. Subsequently, a variety of NeRF-based 3D style transfer methods (Huang et al., 2022; Pang et al., 2023; Liu et al., 2023; Zhang et al., 2024b; Jung et al., 2024) have emerged. Recently, the emergence of 3DGS (Kerbl et al., 2023) has brought new possibilities for scene stylization. It features fast training speed, high rendering quality, and efficient performance. Some works (Liu et al., 2024; Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025) have already explored the potential of stylization under this representation. Currently, 3DGS-based scene stylization methods can be categorized into forward-feed approaches (Liu et al., 2024) and optimization-based approaches (Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025). Forward-feed methods enable zero-shot style transfer. For instance, StyleGaussian (Liu et al., 2024) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (Huang and Belongie, 2017) to achieve real-time stylization. On the other hand, optimization-based methods attain higher-quality style transfer results. Some works (Zhang et al., 2024a; Liu et al., 2025) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
However, existing 3D stylization methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images. Our work focuses on texture transfer. By introducing geometric information, we achieve high-quality results that align with human visual perception.
3. Preliminaries
3.1. 3D Gaussian Splatting
3D Gaussian Splatting (Kerbl et al., 2023) is a novel explicit 3D representation method. It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes. The paramaters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, a opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$ which is represented as spherical harmonic coefficients. Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process. The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$. The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (Lassner and Zollhofer, 2021). Specifically, the Gaussians are grouped into tiles and sorted by depth. The color $C$ of a pixel is computed through $\alpha$-blending:
(1) $C=\sum_{i\in{N}}T_{i}\alpha_{i}c_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$
where $T_{i}$ is the transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian. The initial Gaussians are initialized from a sparse point cloud obtained via Structure-from-Motion (SfM) (Schonberger and Frahm, 2016) using the input multi-view images. During optimization, 3DGS adopts an adaptive densification strategy to control the distribution of Gaussians. Specifically, this strategy performs cloning or splitting of Gaussians based on their positional gradients and sizes. Moreover, it periodically removes Gaussians with low opacity or large geometric extent.
3.2. Optimization for 3D stylization
Optimization for 3D stylization requires balancing both style transfer effects and the preservation of original content information. By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized. Existing optimization-based 3D style transfer methods (Zhang et al., 2022; Pang et al., 2023; Zhang et al., 2024b, a) typically adopt the nearest neighbor feature matching (NNFM) loss (Zhang et al., 2022) as their style loss function. It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them. Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$. The same feature extractor (e.g., VGG (Simonyan and Zisserman, 2014)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively. Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
(2) $L_{style}=\frac{1}{N}\sum_{i,j}\min_{i^{\prime},j^{\prime}}dist(F_{r}(i,j),F_{s}(i^{\prime},j^{\prime})),$
where $N$ is the number of pixels in $F_{r}$, $dist(a,b)$ is the cosine distance between two feature vectors $a$ and $b$.
The content loss $L_{content}$ is the perceptual loss, which is computed using the $L_{2}$ loss between the rendered feature map $F_{r}$ and the content feature map $F_{c}$. The total stylization loss function consists of the style loss and the content loss, expressed as
(3) $L=L_{style}+\lambda L_{content},$
where $\lambda$ is the scaling hyperparameter. By adjusting the scaling hyperparameter $\lambda$, the degree of stylization can be controlled.
4. Method
4.1. GT2-GS FrameWork Overview
In this section, we provide a detailed overview of the proposed GT2-GS framework. During the preprocessing stage, we extract a set of texture features from the input texture image using the proposed Geometry-aware Texture Augmentation (GTA) module. Moreover, Geometry-consistent Texture Loss (GT Loss) is proposed to enable controllable texture transfer. It is computed using the texture features extracted by the GTA module and the prior information from the previous viewpoints. For the input Gaussian, an additional color parameter $c^{g}$ is embedded and initialized using the current color parameter $c$. During 3D texture transfer, we proposed Geometry Preservation Strategy (GPS), which updates Gaussian parameters through multiple iterative rounds. In each round, the color parameter is first switched to $c$, and a texture transfer stage is performed. After Gaussian densification, a geometry correction stage is conducted with the color parameter switched to $c^{g}$. After several rounds of iteration, the texture features are effectively transferred to the scene representation. The pipeline is shown in Fig.  2.
4.2. Geometry-aware Texture Augmentation
The scene learns texture features through rendered images, and the scale of the texture on the object’s surface in the rendered view is related to the distance between the object and the camera. This relationship adheres to the perspective geometry principle. Simultaneously, multi-view rendered images contain rich 2D geometric information such as edges, contours, and shapes. However, texture images exhibit a high degree of repetitive patterns. The texture features extracted using the VGG feature extractor exhibit a homogenization phenomenon. The rendered feature maps fail to align with the desired texture features, thereby impairing the scene’s ability to learn the target texture appearance. To address these issues, we augment the texture features by incorporating geometric information.
To obtain multi-scale texture features that align with the perspective geometric effect, we perform texture mapping on the initial texture image by incorporating the depth map rendered from the original scene. The depth map can be obtained using a method similar to rendering color. Specifically, the depth value $D$ of a pixel is computed by blending $N_{p}$ ordered Gaussians that overlap the pixel as
(4) $D=\sum_{i\in{N_{p}}}T_{i}\alpha_{i}d_{i},$
where $d_{i}$ is the depth for $i$-th Gaussian. Subsequently, the texture image can be scaled according to the depth values. Given the excessive number of distinct depth values for all pixels in the depth map, we sort the values and discretize them into $K$ groups based on the depth intervals. Using the lowest group depth value $Z_{1}$ as the reference, the corresponding scaling factor $s$ for each group is obtained through $s_{k}=Z_{1}/Z_{k}$. The depth value $Z_{1}$ of the lowest group is determined by the upper bound of the depth interval to avoid division-by-zero issues. Meanwhile, the depth values $Z_{k}$ of the remaining groups are determined by the lower bound of their respective intervals. These scaling factors are then converted into scaling matrices $M_{s}$, which are utilized for multi-scale texture mapping by equation as $I_{scale}=M_{s}I_{s}$.
Edges, contours, and shapes, as 2D geometric features, inherently possess directional attributes. The directional information of the entire rendered image can encompass a wide range of angles. However, homogeneous texture images are insufficient to represent geometric information across all directions, particularly for textures such as lines and arrows. To address this, we apply the rotation matrix $M_{r}$ to extend the multi-scale texture in various directions $\Phi$. The complete geometric-aware texture augmentation process can be formulated as
(5) $I_{aug}^{k,\theta}={M_{r}^{\theta}}{M_{s}^{k}}I_{s},$
where $I_{aug}$ is the image obtained through augmentation, $k\in K$ is the index of depth group, and $\theta\in\Phi$ represents the rotation angle applied to the texture image.
Through the aforementioned steps, we have obtained a series of texture images based on geometric information. Although the texture feature set is extensive for scene texture transfer, the excessive number of features within the set adversely impacts the computational efficiency of the texture transfer process. Given the high homogeneity of texture features, we apply center cropping to the extracted feature maps to reduce redundant texture features. Each feature in the texture feature set is defined as $\{f_{k,\theta},s_{k},\theta\}$. The scaling and rotation parameters are retained for subsequent computations.
Figure 3. Differences in perspective. The same textured region (highlighted in red) on the 3D object exhibits different texture orientations under varying viewpoints.
4.3. Geometry-consistent Texture Loss
With the GTA module, we have already obtained sufficient texture features. However, for initial texture images with the same texture patterns but differing directions, the texture feature sets derived from geometry-aware texture enhancement are approximately identical. If the previous NNFM loss is directly used for texture transfer, it becomes impossible to control the texture orientation learned by the scene. To address the issue of controllability, we incorporate the 3D geometric relationships between different camera viewpoints into the calculation of the loss function.
GT Loss is proposed to enable geometry-consistent texture transfer. Specifically, we construct a per-pixel matched target feature map $F_{t}$ for the rendered feature map $F_{r}$. Texture transfer is achieved by minimizing the cosine similarity between the corresponding feature vectors of $F_{r}$ and $F_{t}$. To specify the texture orientation for a given viewpoint, we enforce geometric consistency through inter-view geometric relationships. Specifically, except for the viewpoint with the designated texture orientation, all other viewpoints consider the results from the previous viewpoint when constructing the target feature map. We achieve prior information transmission by using geometric information from the depth map $I_{d}^{v}$ of the current view and intrinsic and extrinsic camera parameters. Through this information, we derive a prior information transmission matrix $M_{p}^{v-1,v}$ via the projection process from the camera pose of current view $\varphi_{v}$ to 3d space to camera pose $\varphi_{v-1}$ in prior view. The computation process of this matrix is formulated as
(6) $M_{p}^{v,v-1}=K_{v-1}[R_{v-1}|T_{v-1}]{[R_{v}|T_{v}]^{-1}}{K_{v}^{-1}},$
where $K_{v}$ and $[R_{v}|T_{v}]$ represent the intrinsic matrix and world-to-camera extrinsic parameters of the $v$-th viewpoint, respectively. Through the information transfer matrix, the current viewpoint can sample the prior feature $f_{k^{\prime},\theta^{\prime}}$ in the screen coordinate system of the prior viewpoint. Considering occlusion relationships, we filter the projected points using the depth map $I_{d}^{v-1}$ of the prior viewpoint. We denote $X$ as a binary mask indicating the positions in the current view’s feature map where prior information can be obtained. Specifically, $X(i,j)=1$ indicates that prior information is available at position $(i,j)$, while $X(i,j)=0$ means it is not.
However, as shown in Fig. 3, the orientation of the same texture region varies across different viewing angles. Each texture feature vector inherently contains scale and orientation. Directly using this feature vector as a prior fails to account for the impact of viewpoint changes in 3D space. To address this, we utilize upsampling to obtain the pixel set $\{p_{v}\}$ corresponding to each feature map location in the pixel coordinate system. Through the transformation relationship $M_{p}^{v-1,v}$, we can determine its corresponding pixel set $\{p_{v-1}\}$ in the previous viewpoint. We use the least squares method to compute the linear transformation matrix $M_{L}\in\mathbb{R}^{2\times 2}$ for texture variation between viewpoints. The obtained transformation matrix $M_{L}$ can be decomposed using SVD to extract the rotation angle $\beta$. The construction method of the texture feature vector at position $(i,j)$ in the target feature map $F_{t}$ is formulated as
(7) $F_{t}(i,j)=\mathop{\arg\min}\limits_{f_{k,\theta},~{}k=k_{(i,j)}}dist(F_{r}(i,j),f_{k,\theta})+\lambda_{p}X(i,j)|\theta^{\prime}+\beta-\theta|,$
where $\lambda_{p}$ is the prior texture orientation control coefficient and $k_{i,j}$ represents the depth group corresponding to the feature position. Subsequently, texture transfer can be achieved by minimizing the cosine similarity between the feature vectors at corresponding positions in the rendered feature map $F_{r}$ and the target feature map $F_{t}$. The geometric-consistent texture transfer loss function is formulated as follows:
(8) $L_{gt}=\frac{1}{N}\sum_{i,j}dist(F_{r}(i,j),F_{t}(i,j)).$
Using only our GT Loss makes it difficult to preserve the original scene content. Therefore, we introduce additional loss functions during the texture transfer stage to retain the scene’s structural information. Needle-like Gaussians negatively affect the quality of texture transfer, making the content difficult to recognize. We measure Gaussian elongation degree using the ratio $l_{max}/l_{mid}$ between the longest axis length $l_{max}$ and the median axis length $l_{mid}$. The regularization term for suppressing needle-like Gaussians is formulated as
(9) $L_{reg}=\frac{1}{N_{G}}\sum_{i}\frac{l_{max}^{i}}{l_{mid}^{i}},$
where $N_{G}$ is the number of Gaussians. The total loss during the texture transfer phase is expressed as
(10) $L_{tot}={\lambda_{gt}}L_{gt}+{\lambda_{c}}L_{content}+{\lambda_{tv}}L_{tv}+{\lambda_{reg}L_{reg}},$
where $L_{tv}$ is the total variation loss, $\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}$ are the coefficients of the corresponding loss functions.
Figure 4. Qualitative Comparison of Texture Transfer.
Figure 5. Qualitative Comparison of Style Transfer.
4.4. Geometry Preservation Strategy
3DGS is an explicit representation, where the scene’s ability to learn textures is closely related to the distribution of Gaussians. In low-texture regions, the density of Gaussians is typically lower. While these regions achieve good rendering quality in the original rendering process, they struggle to learn new texture appearances. The Gaussian densification strategy increases the number of Gaussians in low-texture regions during the texture migration phase, enabling the learning of complex textures. However, Gaussians are densified solely based on gradients. Since texture transfer lacks ground truth during optimization, the densification strategy may introduce erroneous Gaussians floating in space.To address this issue, we propose a two-stage iterative optimization strategy.
Specifically, We perform $N_{iter}$ iterations of optimization. In each iteration, our texture transfer framework first performs the texture transfer stage, followed by the geometry correction stage. During the texture transfer stage, the Gaussian color parameters are set to $c$. We optimize the Gaussian parameters using the total loss function $L_{tot}$ defined in Eq. 10. For the GT Loss $L_{gt}$, we dynamically control the injection of prior information. During the iterative optimization process, prior information $\theta^{\prime}$ and $\beta$ in Eq. 7 are introduced in odd-numbered iterations, while it is omitted in even-numbered iterations. This strategy prevents overly strong priors from adversely affecting fine-grained geometric structures in the scene, such as staircases and railings.
Subsequently, we perform the geometry correction stage to address geometric errors introduced during the densification process. At this stage, the Gaussian parameters are set to the additionally embedded color parameters $c^{g}$. This optimization process employs the original 3DGS reconstruction loss function,
(11) $\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$
where $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ are calculated between the rendered images and the content images. Through an optimization process with ground truth, the Gaussians in the scene are moved to their correct geometric positions. Through multiple iterations of the aforementioned two-stage optimization process, we achieve a balance between texture transfer and geometry preservation.
Figure 6. Qualitative Experiments on Texture Orientation Control.
5. Experiment
Datasets. For the scene datasets, we utilize the LLFF dataset  (Mildenhall et al., 2019) and the T&T dataset (Knapitsch et al., 2017), which are collected from the real world. Additionally, we use images from the ARF (Zhang et al., 2022) style dataset and the DTD dataset (Cimpoi et al., 2014) as reference image datasets.
Baseline. We compare our method with the state-of-the-art 3D stylization method, including ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024). Specifically, ARF is based on NeRF, and ABC-GS and StyleGaussian are based on 3DGS. Optimization-based methods include ARF and ABC-GS, while StyleGaussian is a feed-forward-based method.
5.1. Implementation Details
We perform view-consistent color transfer (Zhang et al., 2022) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters. For the VGG  (Simonyan and Zisserman, 2014) feature extractor, we employ the conv3 block of VGG-16. For texture feature, the GTA module defaults to a depth group count $K$ to 4, and the rotation angle set $\Phi$ covers 360 degrees. Following the GPS, we set the number of iterations $N_{iter}$ to 10 for texture transfer. During the texture transfer stage, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}. Some hyperparameters for forward-facing scenes and 360-degree outdoor scenes differ, and we will provide a more detailed explanation in the Appendix. To ensure a fair comparison with other methods, we use the texture direction of the texture image itself as the prior direction during training. All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
5.2. Qualitative Evaluation
To comprehensively evaluate the effectiveness of our method, the qualitative experiments are divided into three parts: texture transfer, style transfer, and texture orientation control. The results are shown in Fig. 4, Fig. 5 and 6, with more qualitative examples provided in the Appendix.
Texture Transfer. Fig. 4 shows the texture transfer results of our method compared with other methods. Visually, our results exhibit higher fidelity to the reference texture and better alignment with human visual perception. For example, in the geometrically complex orchids scene (Row 1 in Fig. 4), our method produces transfer results that appear as if the texture is naturally wrapped around the surface of the orchids. StyleGaussian, as a zero-shot style transfer method, struggles to handle such complex texture transfers. The results of ABC-GS and ARF appear blurry and exhibit texture discontinuities. These methods show suboptimal texture style matching and visual quality. This is primarily due to the use of identical texture features across multiple views during optimization, which contradicts the geometric consistency of the 3D scene. In comparison, our method leverages the GTA module to extract geometry-aware texture features and applies GT Loss for geometry-consistent appearance optimization.
Style Transfer. To validate the generalizability of our method, we selected artistic style images as reference inputs and conducted a qualitative comparison of style transfer with other methods. As shown in Fig. 5, our approach is more effective in transferring the texture elements of the style image to the 3D scene. In contrast, the results of other methods rely more on rendered views and fail to account for the underlying scene geometry. For instance, in the fortress scene shown in the first row of Figure X, ABC-GS and ARF tend to preserve the color difference between tabletop objects and the table during style transfer. In comparison, our method takes full geometric information into account, allowing textures to be accurately transferred onto object surfaces.
Texture Orientation Control. The proposed GT Loss enables the scene to propagate the target texture information based on the prior view. Therefore, by controlling the learnable texture direction in the initial prior view, we can influence the overall appearance learned by the entire scene. This control is achieved by injecting pseudo prior angle information into the first prior view. As shown in Fig. 6, by combining the reference texture image and view direction control, we achieve high-quality texture transfer results with controllable texture orientation.
Table 1. Quantitative Experiment on Multi-view Consistency.
Methods Short-term Consistency Long-term Consistency
LPIPS($\downarrow$) RMSE($\downarrow$) LPIPS($\downarrow$) RMSE($\downarrow$)
StyleGaussian 0.052 0.051 0.137 0.125
ARF 0.078 0.067 0.175 0.153
ABC-GS 0.051 0.045 0.127 0.118
Ours 0.052 0.045 0.128 0.118
Table 2. Quantitative Experiment on Content Protection.
Ours ABC-GS ARF StyleGaussian
SSIM($\uparrow$) 0.47 0.55 0.34 0.40
CLIP-score($\uparrow$) 0.45 0.39 0.38 0.35
5.3. Quantitative Evaluation
In the 3D scene appearance editing task, maintaining multi-view consistency and preserving scene content information are crucial. To this end, we conducted extensive quantitative experiments from both aspects. We randomly selected 50 pairs of scene-reference image pairs from the scene and reference image datasets to quantitatively evaluate our proposed method.
Multi-view Consistency. We evaluate multi-view consistency using both short-term and long-term consistency metrics. These metrics are computed by warping the rendered image from one viewpoint to another using optical flow (Teed and Deng, 2020). The RMSE and LPIPS scores are then calculated between the warped image and the target view. The results are shown in Table. 1. It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency. Notably, ABC-GS disables the Gaussian densification strategy during optimization. In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPS.
Content Preservation. For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance. We use SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021) to assess the preservation of content information. Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level. In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images. As shown in Table. 4, our proposed method achieves significantly higher scores on both evaluation metrics. In particular, our method outperforms previous approaches in terms of the CLIP-score. It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
Table 3. The results of user study.
Ours ABC-GS ARF StyleGaussian
Texture Alignment($\uparrow$) 4.23 3.02 3.08 2.15
Visual Quality($\uparrow$) 3.90 3.37 3.18 2.22
Figure 7. Ablation study for GT Loss.
Figure 8. Ablation study for GTA and GPS.
5.4. User Study
Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment. Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality. We collected X responses from an online questionnaire distributed through social media platforms. The questionnaire consisted of 20 groups of images along with corresponding prompts. Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint. Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest). As shown in Table. 3, our method consistently outperforms the baseline approaches. Specifically, by comparing the differences across various evaluation dimension, it is evident that users showed a clear preference for our proposed method. This indicates that our approach achieves texture transfer results that better align with human visual perception.
5.5. Ablation Study
Impact of GT Loss. GT Loss is proposed to control the texture orientation. In the ablation study, we perform central rotation on the same texture image to generate multiple inputs with different rotation angles. As shown in Fig. 7, removing the GT Loss leads to identical texture transfer results across different input angles. The reason for this phenomenon is that texture images with different rotation angles produce identical texture feature sets after being processed by the GTA module. Without the constraint of GT Loss, the scene directly learns appearance from the aggregated texture feature set, ignoring orientation variations.
Impact of GTA. As shown in Fig. 8, removing the GTA module results in noticeable texture discontinuities and blurring in the texture transfer outputs. This degradation is attributed to the absence of geometry-aware texture features. Without the GTA module, a large portion of texture features aligned with the scene’s geometry are eliminated. The remaining texture features are insufficient to support learning across the full range of geometric structures within the scene. Learning texture features of the same orientation across different viewpoints fails to achieve continuous texture transfer results between views.
Impact of GPS. The 3D texture transfer tasks inherently lack ground truth supervision, which can introduce incorrect geometry during the scene appearance editing process. This issue is particularly pronounced in the 3DGS-based methods, where appearance and geometry are tightly entangled. To address this, we propose GPS, which leverages the original content images to optimize the geometric parameters of Gaussians and correct inaccurate geometry. As illustrated in Fig. 8, removing GPS results in noticeable artifacts across the scene, especially in challenging 360-degree outdoor environments. This highlights the critical role of GPS in maintaining geometric consistency. Through GPS, our texture transfer framework achieves a balance between learning texture features and preserving scene geometry.
6. Conclusion
In this paper, we introduced GT2-GS, a novel Geometry-aware Texture Transfer framework for Gaussian Splatting that achieves high-quality texture transfer results. Unlike previous 3d stylization methods, our approach explicitly considered the intrinsic relationship between geometry and texture. We observed that feature vectors directly extracted from texture images are insufficient to capture the rich geometric information present in 3D scenes. To address this, we proposed the GTA module, which enhanced the texture feature set by incorporating geometric cues such as perspective geometry, edges, and shapes. Furthermore, we designed a GT Loss that leveraged 3D geometric information across viewpoints to enable controllable texture transfer. To preserve accurate geometry during texture learning, we introduced the GPS strategy to maintain structural integrity throughout the optimization process. Extensive quantitative and qualitative experiments demonstrated the effectiveness of our proposed method. Moreover, our framework was capable of generating high-quality stylization results, showcasing its generalizability.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62472178, 62376244), and Shanghai Urban Digital Transformation Special Fund Project (202301027). We also thank the financial support from Shanghai China Fortune Co., Ltd.
Appendix A ADDITIONAL ANALYSIS
Quantitative Ablation Study on GTA and GPS. We perform a quantitative ablation study on the Geometry-aware Texture Augmentation (GTA) module and the Geometry Preservation Strategy (GPS). The impact of texture transfer on the original scene content is assessed using SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021). We randomly selected 30 pairs of scenes and texture images for the experiment. As shown in Table  4, removing either of these components adversely affects the geometric structure of the scene. Consequently, the scene content becomes less distinguishable, leading to a decrease in semantic similarity with the original scene.
Component-level Impact Analysis of GTA. We conduct a component level ablation study on the proposed GTA module, which includes multi-scale texture enhancement to simulate perspective geometry and multi-angle enhancement to match geometric cues such as edges and contours in the rendered images. As shown in Fig. 9, the multi-angle enhancement significantly impacts the final texture transfer results, while removing the multi-scale enhancement leads to a lack of depth perception in the scene.
Component-level Impact Analysis of GPS. GPS consists of two key components: dynamic prior injection and the geometry correction stage. We conduct ablation studies to investigate their impact on scene geometry. As shown in Fig. 10, removing dynamic prior injection disrupts the geometric structure of textures that are difficult to learn in the specified orientation, resulting in blurry renderings. On the other hand, removing the geometry refinement stage leads to severe Gaussian floating artifacts throughout the scene, severely degrading the 3D geometric integrity.
Impact of Content Loss. As shown in Fig. 11, the content loss controls the strength of texture transfer, helping preserve the original scene content. However, an excessively large $\lambda_{c}$ hinders the scene from effectively learning the desired texture appearance.
Impact of Color Transfer. We follow ARF (Zhang et al., 2022) to perform view-consistent color transfer. As shown in Fig. 12, without color transfer, the scene struggles to learn colors consistent with the texture image, which significantly degrades the visual quality of 3D texture transfer.
Appendix B MORE IMPLEMENTATION DETAILS
For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings. In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively. During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1} and perform $N_{iter}=10$ rounds of GPS optimization, where each round consists of 80 texture transfer steps followed by 400 geometry correction steps. For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005 and use $N_{iter}=6$ rounds with 400 texture transfer steps and 1000 geometry correction steps per round to accommodate more viewpoints.
Figure 9. Component-level Ablation Study of GTA.
Table 4. Quantitative Experiment on Content Protection.
Ours w/o GTA w/o GPS
SSIM($\uparrow$) 0.43 0.37 0.39
CLIP-score($\uparrow$) 0.36 0.35 0.34
Figure 10. Component-level Ablation Study of GPS.
Figure 11. Impact of Content Loss.
Figure 12. Impact of Color Transfer.
Figure 13. Qualitative Comparison of Texture Transfer on the LLFF Dataset.
Figure 14. Qualitative Comparison of Style Transfer on the LLFF Dataset.
Figure 15. Qualitative Comparison of Texture Transfer on the T&T Dataset.
Figure 16. Qualitative Comparison of Style Transfer on the T&T Dataset.
Appendix C MORE QUALITATIVE EVALUATION
For more qualitative evaluation, we compare our method with ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024). Fig. 13, Fig. 14, Fig. 15, and Fig. 16 present the qualitative results.
Appendix D DISCUSSION, LIMITATION, AND FUTURE WORK
Our method primarily consists of three key components: Geometry-aware Texture Augmentation (GTA), Geometric-consistent Texture (GT) Loss, and Geometry Preservation Strategy (GPS). Among them, the proposed GTA module and GT Loss leverage 2D and 3D geometric information respectively, and their computational process is independent of specific 3D representations. Therefore, they can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel. Meanwhile, GPS provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision. However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal. Due to the introduction of a geometry correction stage, our method introduces additional optimization time. Nevertheless, for forward-facing scenes, our method is able to achieve high-quality texture transfer results within 5 minutes.
In future work, we plan to focus on two main aspects. First, we aim to enhance the geometry correction stage of GPS by integrating more accurate 3DGS optimization techniques for geometry reconstruction. This improves the quality of scene appearance editing. In addition, we will explore the integration of diffusion models and multimodal large language models (MLLM) with 3D texture and style transfer. Leveraging the rich prior knowledge of multimodal models, we envision a unified framework that supports both text-driven and image-driven style transfer in 3D scenes.

### v3

@@
 GT2-GS: Geometry-aware Texture Transfer for Gaussian Splatting
-Abstract.
-Transferring 2D textures to 3D modalities is of great significance for improving the efficiency of multimedia content creation.
-Existing approaches have rarely focused on transferring image textures onto 3D representations.
-3D style transfer methods are capable of transferring abstract artistic styles to 3D scenes.
-However, these methods often overlook the geometric information of the scene, which makes it challenging to achieve high-quality 3D texture transfer results.
-In this paper, we present GT2-GS, a geometry-aware texture transfer framework for gaussian splitting.
-From the perspective of matching texture features with geometric information in rendered views, we identify the issue of insufficient texture features and propose a geometry-aware texture augmentation module to expand the texture feature set.
-Moreover, a geometry-consistent texture loss is proposed to optimize texture features into the scene representation.
-This loss function incorporates both camera pose and 3D geometric information of the scene, enabling controllable texture-oriented appearance editing.
-Finally, a geometry preservation strategy is introduced.
-By alternating between the texture transfer and geometry correction stages over multiple iterations, this strategy achieves a balance between learning texture features and preserving geometric integrity.
+Abstract
+Transferring 2D textures onto complex 3D scenes plays a vital role in enhancing the efficiency and controllability of 3D multimedia content creation.
+However, existing 3D style transfer methods primarily focus on transferring abstract artistic styles to 3D scenes.
+These methods often overlook the geometric information of the scene, which makes it challenging to achieve high-quality 3D texture transfer results.
+In this paper, we present GT2-GS, a geometry-aware texture transfer framework for gaussian splatting.
+First, we propose a geometry-aware texture transfer loss that enables view-consistent texture transfer by leveraging prior view-dependent feature information and texture features augmented with additional geometric parameters.
+Moreover, an adaptive fine-grained control module is proposed to address the degradation of scene information caused by low-granularity texture features.
+Finally, a geometry preservation branch is introduced.
+This branch refines the geometric parameters using additionally bound Gaussian color priors, thereby decoupling the optimization objectives of appearance and geometry.
 Extensive experiments demonstrate the effectiveness and controllability of our method.
 Through geometric awareness, our approach achieves texture transfer results that better align with human visual perception.
-Our homepage is available at this https url.
-Texture Transfer, 3D Gaussian Splatting, Geometry-aware
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
-Figure 1.
-Previous methods struggle to accurately transfer complex texture appearances to the scene.
-In contrast, our GT2-GS framework injects geometric information into the scene during texture transfer, enabling geometry-aware texture transfer.
-Our method enables continuous and controllable texture transfer, and is capable of producing perspective effects as illustrated by the blue boxes.
-1. Introduction
-Texture transfer (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) aims to transfer the texture features of a reference image to a target image while preserving the structural and semantic content.
-Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis and were mainly applied to tasks such as image inpainting and data augmentation.
-With the advancement of deep learning, a variety of network-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Cao et al., 2023; Pu et al., 2024) have emerged.
-These methods enable texture transfer by incorporating semantic information, allowing for artistic content creation.
-With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has grown significantly.
-Texture is an essential component of various graphics pipelines and plays a crucial role in producing high-quality 3D assets.
-Traditionally, texture creation requires skilled artists to design them manually, which is both costly and inefficient.
-Recently, the emergence of neural radiance fields (NeRF) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has greatly improved the efficiency of 3D content creation.
-Building upon such 3D representations, it is promising to develop 3D texture transfer method that directly transfers texture features from reference images to 3D representations.
-This approach bypasses the labor-intensive UV mapping process, significantly reducing both the cost and technical barriers of 3D content creation.
-Due to the challenge of multi-view consistency, 2D texture transfer methods (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) cannot be directly applied to 3D scenes.
-Some diffusion-based approaches (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) generate high-quality texture maps for 3D meshes.
-However, these methods modify the appearance of 3D objects via texture maps rather than directly transferring texture features into 3D representations.
-Moreover, they typically operate at the object level, making it difficult to generate textures for entire scenes.
-Therefore, how to achieve efficient 3D texture transfer remains an open and worthwhile research question.
-Existing 3D scene stylization methods (Zhang et al., 2022; Liu et al., 2023, 2024; Zhang et al., 2024a; Liu et al., 2025) focus on transferring the artistic styles of abstract reference images to 3D scenes.
-We attempt to use texture images as references and apply state-of-the-art optimization-based stylization methods to transfer their texture features to 3D scenes.
-However, as shown in the first row on the left side of Fig.
-1, the texture transfer results of this approach are suboptimal.
-Through analysis, we identify the core reason why existing 3D stylization methods cannot be directly applied to texture transfer.
-Texture is closely tied to geometry.
-3D texture transfer needs to be harmonious with both the 3D geometry and the texture features of the reference image, requiring consistency between texture and geometry to ensure strict multi-view consistency.
-However, existing 3D stylization methods overlook the geometry of the scene.
-Firstly, 3D scenes are optimized via gradient descent using losses computed between rendered feature maps and texture features.
-Since texture images typically contain only a single type of texture, applying highly repetitive texture features across diverse rendered views is suboptimal.
-This fails to align with the rich geometric information in the rendered images, such as perspective geometry, edges, contours, and shapes.
-Moreover, current style transfer loss functions are computed independently across views, without considering the 3D geometric relationships between different viewpoints.
-This limits the ability to achieve spatially consistent and controllable texture transfer.
+Our homepage is available at https://vpx-ecnu.github.io/GT2-GS-website.
+Introduction
+3D style transfer (zhang2022arf; zhang2024stylizedgs) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information.
+With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has increased significantly.
+3D stylization techniques offer a promising solution by accelerating the creation of 3D content, particularly in complex 3D scene environments.
+Recently, the emergence of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl3Dgaussians) has significantly advanced the field of 3D stylization.
+NeRF-based stylization methods (zhang2022arf; zhang2023ref; nguyen2022snerf), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry.
+In contrast, 3DGS-based approaches (galerne2025sgsst; liu2025abc; liu2024stylegaussian) offer benefits such as explicit editability and real-time rendering.
+3D stylization methods can be broadly categorized into feed-forward and optimization-based approaches, depending on whether they support zero-shot style transfer.
+While feed-forward methods (liu2023stylerf; huang2022stylizednerf; liu2024stylegaussian) enable zero-shot stylization, they often exhibit lower rendering quality.
+In contrast, optimization-based methods (zhang2022arf; liu2025abc; zhang2024stylizedgs; galerne2025sgsst) typically produce higher-fidelity stylization results.
+Despite these differences, both types of methods commonly define optimization objectives within the VGG feature space during training.
+Figure 1: Previous state-of-the-art methods struggle to accurately transfer complex textures to 3D scene.
+In contrast, our GT2-GS framework incorporates geometric information into the optimization process, enabling geometry-aware texture transfer.
+Texture is a crucial component of various graphics pipelines and plays a vital role in producing high-quality 3D assets.
+Compared to abstract artistic styles, texture elements are more controllable for users.
+However, as illustrated in Fig.
+1, existing methods struggle to accurately transfer the texture of the reference image onto the 3D scene.
+We analyze the limitations of current approaches from the perspective of the scene optimization process.
+First, during scene optimization, multi-view objectives should preserve correct geometric relationships.
+For example, using content images captured from the same scene as ground truth.
+However, existing style transfer methods define their objectives independently across views, without considering the rich geometric structure within the scene or the geometric consistency across viewpoints.
+Moreover, we observe a fine-grained mismatch between the VGG feature space and the pixel space.
+Commonly used style transfer losses fail to account for this mismatch, which leads to regions with high pixel-level information density being easily disrupted by coarse-grained feature representations.
+Addressing these issues is crucial for achieving high-quality texture transfer.
 In this paper, we propose a novel framework GT2-GS for achieving geometry-aware texture transfer.
-Our insight is to integrate comprehensive geometric information into the texture transfer process to achieve controllable results that align with human visual perception.
-To enable geometry-aware texture transfer, we first propose a geometry-aware texture augmentation module.
-This module enhances texture features through scaling and rotation operations based on geometric information from the rendered image, such as perspective, edges, and contours.
-In addtion, a geometry-consistent texture Loss is proposed for 3D texture transfer.
-This loss uses the camera poses and depth geometry between a prior view and the current optimization view, achieving scene appearance editing with controllable texture orientation.
-To balance texture learning and geometry preservation, we further introduce a Geometry Preservation Strategy.
-This strategy alternates the optimization of Gaussian parameters between a texture transfer stage and a geometry correction stage, with multiple iterations performed.
-In this process, the geometry correction stage utilizes content images with ground truth for optimization, ensuring the accurate refinement of scene geometry.
+Our method explicitly accounts for geometric information and fine-grained discrepancies, enabling texture transfer results that better align with human visual perception.
+We first propose a geometry-aware texture transfer loss, which is built upon texture features augmented with additional geometric parameters and incorporates cross-view geometric information to ensure consistent and controllable texture transfer.
+In addition, an adaptive fine-grained control module is proposed.
+It adaptively adjusts the strength of texture learning based on the information density of different pixel regions, thereby mitigating the fine-grained discrepancy between texture features and pixel-level representations.
+To address the coupling between Gaussian geometry and color parameters, we further introduce a geometry preservation branch.
+We bind additional color parameters to the Gaussians and optimize them using the content image as ground truth, in order to accurately refine the scene geometry.
 Extensive experiments demonstrate that our proposed framework achieves high-quality texture transfer results.
 Our main innovations are as follows:
-  • To the best of our knowledge, we are the first to incorporate geometric information into 3D texture transfer.
-The proposed framework enables direction-controllable and high-quality texture transfer results.
-  • Geometry-aware texture Augmentation module and geometry-consistent texture loss are proposed to incorporate the geometric information of the scene during the 3d scene texture transfer.
-Moreover, we introduced a geometric preservation strategy to maintain the correct geometric information of the scene.
+  • We propose a geometry-aware texture transfer framework for general 3DGS scenes. The proposed framework enables controllable and high-quality texture transfer results.
+  • Our Geometry-aware Texture Transfer Loss and Adaptive Fine-Grained Control Module respectively account for geometric information and fine-grained discrepancies between texture features and pixels, enabling high-quality texture transfer.
+Moreover, the Geometry Preservation Branch provides a novel approach to preserving geometry during appearance editing.
   • Extensive experiments demonstrate that, compared to SOTA methods, our proposed work achieves texture transfer results more consistent with human visual perception.
@@
 The source code will be released.
-2. Related Work
-2.1. Texture Transfer
-Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis.
-While preserving the structural information of the target image to achieve texture transfer effects.
+Related Work
+Texture Transfer
+Early texture transfer algorithms (ashikhmin2003fast; lee2010directional; hertzmann2023image; efros2023image) were based on non-parametric texture synthesis.
+These methods preserve the structural information of the target image to achieve texture transfer effects.
 Gatys et al.
-(Gatys et al., 2016) pioneered the use of convolutional neural networks (CNNs) (Simonyan and Zisserman, 2014) for image style transfer, achieving impressive results.
-Subsequently, a variety of deep learning-based style transfer methods (Chen and Schmidt, 2016; Li and Wand, 2016; Huang and Belongie, 2017; Huo et al., 2021) have been proposed.
+(gatys2016image) pioneered the use of convolutional neural networks (CNNs) (simonyan2014very) for image style transfer, achieving impressive results.
+Subsequently, a variety of deep learning-based style transfer methods (chen2016fast; li2016combining; huang2017arbitrary; huo2021manifold) have been proposed.
 These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer.
-Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (Vaswani et al., 2017), some deep learning-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Pu et al., 2024) can transfer local textures from the reference image to the target image based on high-level semantic correspondence.
-With the development of generative models (Goodfellow et al., 2014; Ho et al., 2020), some methods (Lyu et al., 2021; Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) have adopted texture images as guidance conditions for texture transfer.
-Notably, high-quality 3D texture maps can be directly generated using diffusion-based methods (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024).
-However, these methods require precise computation on existing meshes and are difficult to extend to complex and large-scale scenes.
+Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (vaswani2017attention), some deep learning-based texture transfer methods (wang2022texture; chen2022auv; pu2024dynamic) can transfer local textures from the reference image to the target image based on high-level semantic correspondence.
 In this work, we explore how to directly transfer texture features into 3D representations.
-Figure 2.
-Overview of GT2-GS.
+Figure 2: Overview of GT2-GS.
 The overall pipeline is illustrated in (a).
-The input to our proposed framework includes the scene Gaussians, a content image, and a texture image.
-The texture image is transformed into a texture feature set through the GTA module, as shown in (b).
-This feature set is then used to construct the target feature map in the GT Loss, illustrated in (c).
-After embedding the parameter $c^{g}$, the Gaussians are optimized using the two-stage iterative strategy GPS, as shown in (d).
-As a result, the texture features are seamlessly integrated into the Gaussian representation.
-2.2. 3D Style Transfer
+The input to our proposed framework includes the scene Gaussians, content images, and a texture image.
+The texture image is transformed into a texture feature set, which is used to construct the target feature map in the GT2 Loss, as shown in (b).
+We extract pixel-wise information from the input and transform it into an adaptive weight matrix to control texture learning, as illustrated in (c).
+After embedding the parameter $c^{g}$, the Gaussians are optimized using two branches.
+The additional geometry preservation branch is shown in (d).
+As a result, the texture features are integrated into the Gaussian representation.
+3D Style Transfer
 3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency.
-Early research in 3D scene style transfer focused on explicit representations such as point clouds (Cao et al., 2020; Huang et al., 2021) ,meshes (Höllein et al., 2022; Yin et al., 2021), and voxels (Guo et al., 2021; Klehm et al., 2014).
+Early research in 3D scene style transfer focused on explicit representations such as point clouds (cao2020psnet; huang2021learning), meshes (hollein2022stylemesh; yin20213dstylenet), and voxels (guo2021volumetric; klehm2014property).
 Zhang et al.
-(Zhang et al., 2022) explored the potential of NeRF in 3D stylization tasks.
+(zhang2022arf) explored the potential of NeRF in 3D stylization tasks.
 By optimizing the scene representation with the proposed 3D style transfer loss, it achieved pleasing visual effects.
-Subsequently, a variety of NeRF-based 3D style transfer methods (Huang et al., 2022; Pang et al., 2023; Liu et al., 2023; Zhang et al., 2024b; Jung et al., 2024) have emerged.
-Recently, the emergence of 3DGS (Kerbl et al., 2023) has brought new possibilities for scene stylization.
+Subsequently, a variety of NeRF-based 3D style transfer methods (huang2022stylizednerf; pang2023locally; liu2023stylerf; zhang2024coarf; jung2024geometry) have emerged.
+Recently, the emergence of 3DGS (kerbl3Dgaussians) has brought new possibilities for scene stylization.
 It features fast training speed, high rendering quality, and efficient performance.
-Some works (Liu et al., 2024; Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025) have already explored the potential of stylization under this representation.
-Currently, 3DGS-based scene stylization methods can be categorized into forward-feed approaches (Liu et al., 2024) and optimization-based approaches (Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025).
-Forward-feed methods enable zero-shot style transfer.
-For instance, StyleGaussian (Liu et al., 2024) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (Huang and Belongie, 2017) to achieve real-time stylization.
+Some works (liu2024stylegaussian; zhang2024stylizedgs; mei2024regs; liu2025abc) have already explored the potential of stylization under this representation.
+Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (liu2024stylegaussian) and optimization-based approaches (zhang2024stylizedgs; mei2024regs; liu2025abc).
+Feed-forward methods enable zero-shot style transfer.
+For instance, StyleGaussian (liu2024stylegaussian) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (huang2017arbitrary) to achieve real-time stylization.
 On the other hand, optimization-based methods attain higher-quality style transfer results.
-Some works (Zhang et al., 2024a; Liu et al., 2025) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
-However, existing 3D stylization methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images.
-Our work focuses on texture transfer.
-By introducing geometric information, we achieve high-quality results that align with human visual perception.
-3. Preliminaries
-3.1. 3D Gaussian Splatting
-3D Gaussian Splatting (Kerbl et al., 2023) is a novel explicit 3D representation method.
+Some works (zhang2024stylizedgs; liu2025abc) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
+However, existing methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images.
+Meanwhile, these methods overlook the fine-grained discrepancies between features and pixels.
+By incorporating geometric information and adaptive fine-grained control, our method is able to achieve high-quality texture transfer results.
+Preliminaries
+3D Gaussian Splatting
+3D Gaussian Splatting (kerbl3Dgaussians) is a novel explicit 3D representation method.
 It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes.
-The paramaters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, a opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$ which is represented as spherical harmonic coefficients.
+The parameters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, an opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$, which is represented as spherical harmonic coefficients.
 Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process.
 The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$.
-The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (Lassner and Zollhofer, 2021).
+The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (lassner2021pulsar).
 Specifically, the Gaussians are grouped into tiles and sorted by depth.
 The color $C$ of a pixel is computed through $\alpha$-blending:
-(1) $C=\sum_{i\in{N}}T_{i}\alpha_{i}c_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$
-where $T_{i}$ is the transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian.
-The initial Gaussians are initialized from a sparse point cloud obtained via Structure-from-Motion (SfM) (Schonberger and Frahm, 2016) using the input multi-view images.
+$C=\sum_{i\in N}T_{i}\alpha_{i}c_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
+where $T_{i}$ is the transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian. The depth value $D$ can be obtained using a method similar to rendering color,
+$D=\sum_{i\in N}T_{i}\alpha_{i}d_{i},$ (2)
+where $d_{i}$ is the depth for $i$-th Gaussian.
 During optimization, 3DGS adopts an adaptive densification strategy to control the distribution of Gaussians.
 Specifically, this strategy performs cloning or splitting of Gaussians based on their positional gradients and sizes.
-Moreover, it periodically removes Gaussians with low opacity or large geometric extent.
-3.2. Optimization for 3D stylization
-Optimization for 3D stylization requires balancing both style transfer effects and the preservation of original content information.
+Style Transfer Loss
 By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized.
-Existing optimization-based 3D style transfer methods (Zhang et al., 2022; Pang et al., 2023; Zhang et al., 2024b, a) typically adopt the nearest neighbor feature matching (NNFM) loss (Zhang et al., 2022) as their style loss function.
+Existing optimization-based 3D style transfer methods (zhang2022arf; pang2023locally; zhang2024coarf; zhang2024stylizedgs) typically adopt the nearest neighbor feature matching (NNFM) loss (zhang2022arf) as their style loss function.
 It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them.
 Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$.
-The same feature extractor (e.g., VGG (Simonyan and Zisserman, 2014)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively.
+The same feature extractor (e.g., VGG (simonyan2014very)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively.
 Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
-(2) $L_{style}=\frac{1}{N}\sum_{i,j}\min_{i^{\prime},j^{\prime}}dist(F_{r}(i,j),F_{s}(i^{\prime},j^{\prime})),$
-where $N$ is the number of pixels in $F_{r}$, $dist(a,b)$ is the cosine distance between two feature vectors $a$ and $b$.
-The content loss $L_{content}$ is the perceptual loss, which is computed using the $L_{2}$ loss between the rendered feature map $F_{r}$ and the content feature map $F_{c}$.
-The total stylization loss function consists of the style loss and the content loss, expressed as
-(3) $L=L_{style}+\lambda L_{content},$
-where $\lambda$ is the scaling hyperparameter. By adjusting the scaling hyperparameter $\lambda$, the degree of stylization can be controlled.
-4. Method
-4.1. GT2-GS FrameWork Overview
+$L_{style}=\frac{1}{N}\sum_{i,j}\min_{i^{\prime},j^{\prime}}dist(F_{r}(i,j),F_{s}(i^{\prime},j^{\prime})),$ (3)
+where $N$ is the number of pixels in $F_{r}$, $dist(a,b)$ is the cosine distance between two feature vectors $a$ and $b$:
+$dist(a,b)=1-\frac{a\cdot b}{\|a\|\|b\|}.$ (4)
+Method
+GT2-GS Framework Overview
 In this section, we provide a detailed overview of the proposed GT2-GS framework.
-During the preprocessing stage, we extract a set of texture features from the input texture image using the proposed Geometry-aware Texture Augmentation (GTA) module.
-Moreover, Geometry-consistent Texture Loss (GT Loss) is proposed to enable controllable texture transfer.
-It is computed using the texture features extracted by the GTA module and the prior information from the previous viewpoints.
-For the input Gaussian, an additional color parameter $c^{g}$ is embedded and initialized using the current color parameter $c$.
-During 3D texture transfer, we proposed Geometry Preservation Strategy (GPS), which updates Gaussian parameters through multiple iterative rounds.
-In each round, the color parameter is first switched to $c$, and a texture transfer stage is performed.
-After Gaussian densification, a geometry correction stage is conducted with the color parameter switched to $c^{g}$.
-After several rounds of iteration, the texture features are effectively transferred to the scene representation.
+Geometry-aware Texture Transfer Loss (GT2 Loss) is proposed to enable controllable and high-quality texture transfer.
+It is computed using texture features bound with geometric parameters and the prior information from the previous viewpoints.
+Furthermore, we propose the Adaptive Fine-grained Control Module (AFCM) to regulate the texture learning intensity based on the scene information density in the pixel space.
+This prevents excessive information from being corrupted by low-granularity features.
+Finally, for the input Gaussian, an additional color parameter $c^{g}$ is embedded and initialized using the current color parameter $c$.
+The Geometry Preservation Branch (GPB) is introduced based on $c^{g}$, with the aim of decoupling the optimization processes of the geometry and appearance of the scene.
 The pipeline is shown in Fig.
 2.
-4.2. Geometry-aware Texture Augmentation
-The scene learns texture features through rendered images, and the scale of the texture on the object’s surface in the rendered view is related to the distance between the object and the camera.
-This relationship adheres to the perspective geometry principle.
-Simultaneously, multi-view rendered images contain rich 2D geometric information such as edges, contours, and shapes.
-However, texture images exhibit a high degree of repetitive patterns.
-The texture features extracted using the VGG feature extractor exhibit a homogenization phenomenon.
-The rendered feature maps fail to align with the desired texture features, thereby impairing the scene’s ability to learn the target texture appearance.
-To address these issues, we augment the texture features by incorporating geometric information.
-To obtain multi-scale texture features that align with the perspective geometric effect, we perform texture mapping on the initial texture image by incorporating the depth map rendered from the original scene.
-The depth map can be obtained using a method similar to rendering color.
-Specifically, the depth value $D$ of a pixel is computed by blending $N_{p}$ ordered Gaussians that overlap the pixel as
-(4) $D=\sum_{i\in{N_{p}}}T_{i}\alpha_{i}d_{i},$
-where $d_{i}$ is the depth for $i$-th Gaussian.
-Subsequently, the texture image can be scaled according to the depth values.
-Given the excessive number of distinct depth values for all pixels in the depth map, we sort the values and discretize them into $K$ groups based on the depth intervals.
-Using the lowest group depth value $Z_{1}$ as the reference, the corresponding scaling factor $s$ for each group is obtained through $s_{k}=Z_{1}/Z_{k}$.
-The depth value $Z_{1}$ of the lowest group is determined by the upper bound of the depth interval to avoid division-by-zero issues.
-Meanwhile, the depth values $Z_{k}$ of the remaining groups are determined by the lower bound of their respective intervals.
-These scaling factors are then converted into scaling matrices $M_{s}$, which are utilized for multi-scale texture mapping by equation as $I_{scale}=M_{s}I_{s}$.
-Edges, contours, and shapes, as 2D geometric features, inherently possess directional attributes.
-The directional information of the entire rendered image can encompass a wide range of angles.
-However, homogeneous texture images are insufficient to represent geometric information across all directions, particularly for textures such as lines and arrows.
-To address this, we apply the rotation matrix $M_{r}$ to extend the multi-scale texture in various directions $\Phi$.
-The complete geometric-aware texture augmentation process can be formulated as
-(5) $I_{aug}^{k,\theta}={M_{r}^{\theta}}{M_{s}^{k}}I_{s},$
-where $I_{aug}$ is the image obtained through augmentation, $k\in K$ is the index of depth group, and $\theta\in\Phi$ represents the rotation angle applied to the texture image.
-Through the aforementioned steps, we have obtained a series of texture images based on geometric information.
-Although the texture feature set is extensive for scene texture transfer, the excessive number of features within the set adversely impacts the computational efficiency of the texture transfer process.
-Given the high homogeneity of texture features, we apply center cropping to the extracted feature maps to reduce redundant texture features.
-Each feature in the texture feature set is defined as $\{f_{k,\theta},s_{k},\theta\}$.
-The scaling and rotation parameters are retained for subsequent computations.
-Figure 3. Differences in perspective. The same textured region (highlighted in red) on the 3D object exhibits different texture orientations under varying viewpoints.
-4.3. Geometry-consistent Texture Loss
-With the GTA module, we have already obtained sufficient texture features.
-However, for initial texture images with the same texture patterns but differing directions, the texture feature sets derived from geometry-aware texture enhancement are approximately identical.
-If the previous NNFM loss is directly used for texture transfer, it becomes impossible to control the texture orientation learned by the scene.
-To address the issue of controllability, we incorporate the 3D geometric relationships between different camera viewpoints into the calculation of the loss function.
-GT Loss is proposed to enable geometry-consistent texture transfer.
-Specifically, we construct a per-pixel matched target feature map $F_{t}$ for the rendered feature map $F_{r}$.
+Figure 3: Differences in Perspective. The same textured region (highlighted in red) on the 3D object exhibits different texture orientations under varying viewpoints.
+Geometry-aware Texture Transfer Loss
+Existing methods primarily focus on artistic style transfer and are mostly based on NNFM loss.
+As shown in Eq.
+3, the objective considers only the relationship between the rendered feature map and the style feature map.
+The optimization objectives across different viewpoints are constructed independently, without accounting for the underlying geometric relationships between views.
+However, texture is inherently tied to the geometry of the scene.
+To achieve accurate texture transfer, we incorporate geometric information into the loss function computation during the optimization process.
+First, we associate geometric parameters with each texture feature.
+Considering the perspective geometry within the scene, edge geometric structures, and transformations across viewpoints, we apply scaling and rotation operations to the texture images.
+Texture features are then extracted from the transformed images and aggregated into a feature set.
+Specifically, we obtain the original scene depth map using Eq.
+2.
+The depth values are then sorted and discretized into $K$ groups based on predefined depth intervals.
+The scaling factor for each group is computed as the ratio between the lowest group depth value $Z_{1}$ and the group’s depth value $Z_{k}$.
+Next, each scaled image is rotated by an angle $\theta$ to capture multi-directional information present in rendered views.
+Each feature in the texture feature set is denoted as $\{f_{k,\theta}\}$.
+The scaling parameter $k$ and rotation angle $\theta$ are retained for subsequent computations.
+Based on the constructed set of texture features, we propose the GT2 Loss to enable geometry-consistent texture transfer.
+We construct a per-pixel matched target feature map $F_{t}$ for the rendered feature map $F_{r}$.
 Texture transfer is achieved by minimizing the cosine similarity between the corresponding feature vectors of $F_{r}$ and $F_{t}$.
-To specify the texture orientation for a given viewpoint, we enforce geometric consistency through inter-view geometric relationships.
-Specifically, except for the viewpoint with the designated texture orientation, all other viewpoints consider the results from the previous viewpoint when constructing the target feature map.
-We achieve prior information transmission by using geometric information from the depth map $I_{d}^{v}$ of the current view and intrinsic and extrinsic camera parameters.
-Through this information, we derive a prior information transmission matrix $M_{p}^{v-1,v}$ via the projection process from the camera pose of current view $\varphi_{v}$ to 3d space to camera pose $\varphi_{v-1}$ in prior view.
-The computation process of this matrix is formulated as
-(6) $M_{p}^{v,v-1}=K_{v-1}[R_{v-1}|T_{v-1}]{[R_{v}|T_{v}]^{-1}}{K_{v}^{-1}},$
-where $K_{v}$ and $[R_{v}|T_{v}]$ represent the intrinsic matrix and world-to-camera extrinsic parameters of the $v$-th viewpoint, respectively.
-Through the information transfer matrix, the current viewpoint can sample the prior feature $f_{k^{\prime},\theta^{\prime}}$ in the screen coordinate system of the prior viewpoint.
+To ensure that the feature maps $F_{t}$ constructed from multiple viewpoints maintain correct geometric relationships, we incorporate the construction result of the previous viewpoint’s feature map $F_{t}^{v-1}$ when building the target feature map $F_{t}^{v}$ for the current viewpoint.
+The computation process of the homography matrix $M_{p}^{v,v-1}$ between two viewpoints is formulated as
+$M_{p}^{v,v-1}=K_{v-1}[R_{v-1}|T_{v-1}]{[R_{v}|T_{v}]^{-1}}{K_{v}^{-1}},$ (5)
+where $K_{v}$ and $R_{v}|T_{v}]$ represent the intrinsic matrix and world-to-camera extrinsic parameters of the $v$-th viewpoint, respectively.
+Through the homography matrix, the current viewpoint can sample the prior feature $f_{k^{\prime},\theta^{\prime}}$ in the screen coordinate system of the prior viewpoint.
 Considering occlusion relationships, we filter the projected points using the depth map $I_{d}^{v-1}$ of the prior viewpoint.
-We denote $X$ as a binary mask indicating the positions in the current view’s feature map where prior information can be obtained.
-Specifically, $X(i,j)=1$ indicates that prior information is available at position $(i,j)$, while $X(i,j)=0$ means it is not.
 However, as shown in Fig.
@@
 The construction method of the texture feature vector at position $(i,j)$ in the target feature map $F_{t}$ is formulated as
-(7) $F_{t}(i,j)=\mathop{\arg\min}\limits_{f_{k,\theta},~{}k=k_{(i,j)}}dist(F_{r}(i,j),f_{k,\theta})+\lambda_{p}X(i,j)|\theta^{\prime}+\beta-\theta|,$
-where $\lambda_{p}$ is the prior texture orientation control coefficient and $k_{i,j}$ represents the depth group corresponding to the feature position.
-Subsequently, texture transfer can be achieved by minimizing the cosine similarity between the feature vectors at corresponding positions in the rendered feature map $F_{r}$ and the target feature map $F_{t}$.
-The geometric-consistent texture transfer loss function is formulated as follows:
-(8) $L_{gt}=\frac{1}{N}\sum_{i,j}dist(F_{r}(i,j),F_{t}(i,j)).$
-Using only our GT Loss makes it difficult to preserve the original scene content.
-Therefore, we introduce additional loss functions during the texture transfer stage to retain the scene’s structural information.
-Needle-like Gaussians negatively affect the quality of texture transfer, making the content difficult to recognize.
-We measure Gaussian elongation degree using the ratio $l_{max}/l_{mid}$ between the longest axis length $l_{max}$ and the median axis length $l_{mid}$.
-The regularization term for suppressing needle-like Gaussians is formulated as
-(9) $L_{reg}=\frac{1}{N_{G}}\sum_{i}\frac{l_{max}^{i}}{l_{mid}^{i}},$
-where $N_{G}$ is the number of Gaussians. The total loss during the texture transfer phase is expressed as
-(10) $L_{tot}={\lambda_{gt}}L_{gt}+{\lambda_{c}}L_{content}+{\lambda_{tv}}L_{tv}+{\lambda_{reg}L_{reg}},$
-where $L_{tv}$ is the total variation loss, $\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}$ are the coefficients of the corresponding loss functions.
-Figure 4. Qualitative Comparison of Texture Transfer.
-Figure 5. Qualitative Comparison of Style Transfer.
-4.4. Geometry Preservation Strategy
+$F_{t}(i,j)=\mathop{\arg\min}\limits_{f_{k,\theta}}dist(F_{r}(i,j),f_{k,\theta})+\lambda_{p}|\theta^{\prime}+\beta-\theta|,$ (6)
+where $\lambda_{p}$ is the prior texture orientation control coefficient.
+Subsequently, texture transfer can be achieved by minimizing the cosine similarity between the feature vectors at corresponding positions in the rendered feature map $F_{r}^{v}$ and the target feature map $F_{t}^{v}$ for the current viewpoint $v$.
+The geometry-aware texture transfer loss function is formulated as follows:
+$L_{gt}=\frac{1}{N}\sum_{i,j}dist(F_{r}^{v}(i,j),F_{t}^{v}(i,j)).$ (7)
+Adaptive Fine-grained Control Module
+Most existing optimization-based style transfer methods rely on VGG features.
+However, feature maps extracted through multiple layers of convolutional neural networks exhibit significantly lower granularity compared to the original image pixels.
+When optimizing the scene using only texture transfer losses, this can lead to the following issues: First, due to perspective projection, regions with greater depth tend to concentrate more scene information.
+Learning coarse-grained texture features may overwrite these important details.
+Second, scenes often contain geometrically fine-grained structures such as stairs and railings, which can be degraded or lost under coarse texture representations.
+To address these challenges, we propose an Adaptive Fine-Grained Control Module.
+Specifically, we first obtain the depth map $I_{d}$ and frequency density map $I_{f}$ for each viewpoint from the original Gaussian scene and content image, respectively.
+Both are rescaled to match the spatial dimensions of the feature maps.
+Furthermore, we introduce a geometric distortion map $\Phi$, defined as the discrepancy in geometric information between the learned texture features and rendered features.
+The geometric distortion map is computed by measuring the angular difference between the features obtained using Eq.
+6 and those derived without any prior information.
+The adaptive fine-grained control map is formulated as
+$W^{v}=\lambda_{d}(1-I_{d}^{v})+\lambda_{f}(1-I_{f}^{v})+\lambda_{\Phi}(1-\Phi),$ (8)
+where $I_{d}^{v}$, $I_{f}^{v}$ and $\Phi$ are all normalized, and $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{\Phi}$ are weight coefficients.
+In most cases, we aim to alter the appearance of foreground objects while simultaneously satisfying requirements for both shallow depth and high frequency.
+We therefore employ an additive formulation to balance the relationship between depth and frequency.
+Furthermore, we regularize the learning process to favor textures with low distortion, thereby better preserving geometric fidelity during texture adaptation.
+The derived adaptive weight matrix $W^{v}$ is subsequently applied to Eq. 7 as
+$L_{wgt}=\frac{1}{N}\sum_{i,j}W^{v}(i,j)\ dist(F_{r}^{v}(i,j),F_{t}^{v}(i,j)),$ (9)
+where $L_{wgt}$ is weighted GT2 Loss. The total loss during the texture transfer phase is expressed as
+$L_{tot}={\lambda_{wgt}}L_{wgt}+{\lambda_{c}}L_{content}+{\lambda_{tv}}L_{tv},$ (10)
+where $L_{tv}$ is the total variation loss, $\lambda_{wgt}$, $\lambda_{c}$, $\lambda_{tv}$ are the coefficients of the corresponding loss functions.
+Figure 4: Comparison of Geometry Preservation Results.
+It can be observed that as the number of Gaussians increases, depth regularization struggles to preserve the geometric information of the scene.
+$\lambda_{dep}$ is the weighting coefficient for the depth regularization term.
+Figure 5: Qualitative Comparison of Texture Transfer and Style Transfer.
+The first four rows in the figure show the texture transfer results, while the last two rows present the style transfer results.
+Geometry Preservation Branch
 3DGS is an explicit representation, where the scene’s ability to learn textures is closely related to the distribution of Gaussians.
@@
 However, Gaussians are densified solely based on gradients.
-Since texture transfer lacks ground truth during optimization, the densification strategy may introduce erroneous Gaussians floating in space.To address this issue, we propose a two-stage iterative optimization strategy.
-Specifically, We perform $N_{iter}$ iterations of optimization.
-In each iteration, our texture transfer framework first performs the texture transfer stage, followed by the geometry correction stage.
-During the texture transfer stage, the Gaussian color parameters are set to $c$.
-We optimize the Gaussian parameters using the total loss function $L_{tot}$ defined in Eq.
-10.
-For the GT Loss $L_{gt}$, we dynamically control the injection of prior information.
-During the iterative optimization process, prior information $\theta^{\prime}$ and $\beta$ in Eq.
-7 are introduced in odd-numbered iterations, while it is omitted in even-numbered iterations.
-This strategy prevents overly strong priors from adversely affecting fine-grained geometric structures in the scene, such as staircases and railings.
-Subsequently, we perform the geometry correction stage to address geometric errors introduced during the densification process.
-At this stage, the Gaussian parameters are set to the additionally embedded color parameters $c^{g}$.
-This optimization process employs the original 3DGS reconstruction loss function,
-(11) $\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$
-where $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ are calculated between the rendered images and the content images.
+Since texture transfer lacks ground truth during optimization, the densification strategy may introduce erroneous Gaussians floating in space.
+As shown in Fig.
+4, simply adding depth regularization does not solve this problem.
+Unlike NeRF, explicit 3D Gaussians jointly encode both geometric and color parameters.
+To address this issue, we propose a specialized branch for optimizing the geometric parameters of Gaussians.
+Our key insight is to introduce an additional optimization objective focused on geometry preservation during training, in order to balance appearance optimization with geometric integrity.
+Specifically, we associate each Gaussian with an additional color parameter $c^{g}$, initialized using the original color values from the scene.
+During optimization, we render an image $I_{g}$ using these color parameters and optimize the Gaussian parameters by treating the content image $I_{c}$ as ground truth.
+The 3D Gaussian Splatting reconstruction loss function is formulated as
+$\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$ (11)
+where $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ are calculated between the rendered image $I_{g}$ and the content image $I_{c}$.
 Through an optimization process with ground truth, the Gaussians in the scene are moved to their correct geometric positions.
-Through multiple iterations of the aforementioned two-stage optimization process, we achieve a balance between texture transfer and geometry preservation.
-Figure 6. Qualitative Experiments on Texture Orientation Control.
-5. Experiment
+Experiment
 Datasets.
-For the scene datasets, we utilize the LLFF dataset  (Mildenhall et al., 2019) and the T&T dataset (Knapitsch et al., 2017), which are collected from the real world.
-Additionally, we use images from the ARF (Zhang et al., 2022) style dataset and the DTD dataset (Cimpoi et al., 2014) as reference image datasets.
+For the scene datasets, we utilize the LLFF dataset (mildenhall2019local) and the Tanks and Temples (T&T) dataset (knapitsch2017tanks), which are collected from the real world.
+Additionally, we use images from the ARF (zhang2022arf) style dataset and the DTD dataset (cimpoi2014describing) as reference image datasets.
 Baseline.
-We compare our method with the state-of-the-art 3D stylization method, including ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024).
-Specifically, ARF is based on NeRF, and ABC-GS and StyleGaussian are based on 3DGS.
-Optimization-based methods include ARF and ABC-GS, while StyleGaussian is a feed-forward-based method.
-5.1. Implementation Details
-We perform view-consistent color transfer (Zhang et al., 2022) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters.
-For the VGG  (Simonyan and Zisserman, 2014) feature extractor, we employ the conv3 block of VGG-16.
-For texture feature, the GTA module defaults to a depth group count $K$ to 4, and the rotation angle set $\Phi$ covers 360 degrees.
-Following the GPS, we set the number of iterations $N_{iter}$ to 10 for texture transfer.
-During the texture transfer stage, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}.
-Some hyperparameters for forward-facing scenes and 360-degree outdoor scenes differ, and we will provide a more detailed explanation in the Appendix.
-To ensure a fair comparison with other methods, we use the texture direction of the texture image itself as the prior direction during training.
+We compare our method with the state-of-the-art 3D stylization methods, including SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref), and SNeRF (nguyen2022snerf).
+Specifically, SGSST, ABC-GS, and StyleGaussian are based on 3DGS, and ARF, Ref-NPR, and SNeRF are based on NeRF.
+StyleGaussian is a feed-forward-based method, while others are optimization-based methods.
+Implementation Details
+We perform view-consistent color transfer (zhang2022arf) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters.
+For the VGG (simonyan2014very) feature extractor, we employ the conv3 block of VGG-16.
+When associating color parameters with texture features, we set the default number of depth groups $K$ to 4, and the rotation angle $\theta$ is sampled across the full 360 degrees.
+The hyperparameter for AFCM is denoted as $\{\lambda_{d},\lambda_{f},\lambda_{\Phi}\}$ = {0.8, 0.8, 0.25}.
+For texture transfer optimization, we set $\{\lambda_{wgt},\lambda_{c},\lambda_{tv}\}$ = {2, 0.005, 0.02}.
 All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
-5.2. Qualitative Evaluation
-To comprehensively evaluate the effectiveness of our method, the qualitative experiments are divided into three parts: texture transfer, style transfer, and texture orientation control.
-The results are shown in Fig.
-4, Fig.
-5 and 6, with more qualitative examples provided in the Appendix.
+Qualitative Evaluation
+To comprehensively evaluate the effectiveness of our method, the qualitative experiments are divided into two parts: texture transfer and style transfer. The results are shown in Fig. 5.
 Texture Transfer.
@@
 For example, in the geometrically complex orchids scene (row 1 in Fig.
-4), our method produces transfer results that appear as if the texture is naturally wrapped around the surface of the orchids.
+5), our method produces transfer results that appear as if the texture is wrapped naturally around the surface of the orchids.
 StyleGaussian, as a zero-shot style transfer method, struggles to handle such complex texture transfers.
-The results of ABC-GS and ARF appear blurry and exhibit texture discontinuities.
-These methods show suboptimal texture style matching and visual quality.
-This is primarily due to the use of identical texture features across multiple views during optimization, which contradicts the geometric consistency of the 3D scene.
-In comparison, our method leverages the GTA module to extract geometry-aware texture features and applies GT Loss for geometry-consistent appearance optimization.
+The results of other optimization-based methods exhibit texture discontinuities and appear noticeably blurry, significantly deteriorating overall visual quality.
+This is primarily due to the inherent coupling between texture and geometry, which these methods fail to consider during optimization.
+In contrast, our geometry-aware approach ensures consistent texture transfer across multiple views and coherent 3D results.
 Style Transfer.
@@
 5, our approach is more effective in transferring the texture elements of the style image to the 3D scene.
-In contrast, the results of other methods rely more on rendered views and fail to account for the underlying scene geometry.
-For instance, in the fortress scene shown in the first row of Figure X, ABC-GS and ARF tend to preserve the color difference between tabletop objects and the table during style transfer.
-In comparison, our method takes full geometric information into account, allowing textures to be accurately transferred onto object surfaces.
+SGSST and ABC-GS disable the Gaussian densification strategy during style transfer optimization, making it challenging to capture rich and detailed style features.
+For instance, in the trex scene (second-to-last row of Fig.
+5), the appearance of the white wall region fails to be effectively optimized.
+In contrast, our method benefits from the joint effect of the AFCM and GPB.
+This design enables the model to preserve scene geometry while effectively learning style textures from the reference image.
+NeRF-based methods tend to capture more abstract stylistic elements, such as brush strokes.
+Methods SSIM($\uparrow$) CLIP-score($\uparrow$) ST-LPIPS($\downarrow$) ST-RMSE($\downarrow$) LT-LPIPS($\downarrow$) LT-RMSE($\downarrow$)
+Ours 0.51 0.47 0.054 0.048 0.087 0.077
+SGSST 0.45 0.44 0.075 0.072 0.119 0.108
+ABC-GS 0.56 0.46 0.049 0.041 0.080 0.068
+StyleGaussian 0.41 0.40 0.058 0.052 0.097 0.082
+ARF 0.37 0.45 0.109 0.072 0.152 0.108
+Ref-NPR 0.35 0.42 0.092 0.069 0.137 0.102
+SNeRF 0.48 0.36 0.075 0.057 0.127 0.090
+Table 1: Quantitative Experiment on Multi-view Consistency and Content Preservation.
+Quantitative Evaluation
+In the 3D scene appearance editing task, maintaining multi-view consistency and preserving scene content information are crucial.
+To this end, we conducted extensive quantitative experiments from both aspects.
+We randomly selected 100 scene–reference image pairs to quantitatively evaluate our method.
+Multi-view Consistency.
+We evaluate multi-view consistency (liu2024stylegaussian) using both short-term and long-term consistency metrics.
+The results are shown in Tab.
+1.
+It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency.
+Notably, ABC-GS disables the Gaussian densification strategy during optimization.
+In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPB.
+Content Preservation.
+For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance.
+We use SSIM (wang2004image) and CLIP-score (radford2021learning) to assess the preservation of content information.
+Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level.
+In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images.
+As shown in Tab.
+1, our proposed method achieves significantly higher scores on both evaluation metrics.
+In particular, our method outperforms previous approaches in terms of the CLIP-score.
+It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
+Figure 6: Ablation study for GT2 Loss, AFCM, and GPB.
+Ablation Study
+Impact of GT2 Loss.
+GT2 Loss is introduced to incorporate geometric information into the optimization objective of texture transfer.
+As shown in Fig.
+6, removing the GT2 Loss results in noticeable texture discontinuities and blurring in the texture transfer outputs.
+Tab.
+2 further demonstrates the significance of GT2 Loss from a quantitative perspective.
+Impact of AFCM.
+The AFCM is introduced to address the granularity mismatch between texture features and pixel-level details.
+It encourages texture learning in foreground and low-frequency regions, while preserving scene content in background and high-frequency areas.
+As illustrated in Fig.
+6, in the fern scene (second row), low-texture foreground regions fail to capture style patterns without the presence of AFCM.
+In the truck scene (third row), which exhibits significant depth variation as a 360° environment, removing AFCM leads to noticeable degradation of geometric and appearance fidelity.
+Impact of GPB.
+The 3D texture transfer tasks inherently lack ground truth supervision, which can introduce incorrect geometry during the scene appearance editing process.
+To address this, we propose GPB, which leverages the original content images to optimize the geometric parameters of Gaussians and correct inaccurate geometry.
+As illustrated in Fig.
+6, removing GPB results in noticeable artifacts across the scene.
+Tab.
+2 highlights the importance of GPB in maintaining content fidelity.
+Through GPB, our texture transfer framework achieves a balance between learning texture features and preserving scene geometry.
+Ours w/o GT2 Loss w/o AFCM w/o GPB
+SSIM($\uparrow$) 0.41 0.38 0.45 0.31
+CLIP-score($\uparrow$) 0.39 0.36 0.38 0.37
+Table 2: Quantitative Ablation Study on Content Protection. The results are obtained from 25 randomly selected experiments conducted on LLFF scenes.
+Conclusion
+In this paper, we introduced GT2-GS, a novel Geometry-aware Texture Transfer framework for Gaussian Splatting that achieves high-quality texture transfer results.
+Unlike the previous 3D stylization methods, our approach explicitly considered the intrinsic relationship between geometry and texture.
+We first propose the GT2 Loss, which leverages features augmented with geometric parameters and utilizes cross-view priors to guide scene optimization, enabling geometrically consistent texture transfer.
+AFCM addresses the granularity mismatch between features and pixels by adaptively controlling the strength of texture learning.
+Additionally, GPB introduces a geometry optimization objective grounded in ground-truth appearance, effectively preserving scene geometry during complex stylization.
+Extensive quantitative and qualitative experiments demonstrated the effectiveness of our proposed method.
+Moreover, our framework was capable of generating high-quality stylization results, showcasing its generalizability.
+As a limitation, because optimizing rendered VGG features requires minimizing texture cosine distance while preserving content loss, which results in a texture interpolated between scene and texture geometry.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (62572194, 62472178, 62376244) and the Fundamental Research Funds for the Central Universities.
+This work was supported by the Shanghai Frontiers Science Center of Molecule Intelligent Syntheses.
+Meanwhile, we thank the financial support from Shanghai Chinafortune Co., Ltd.
+ADDITIONAL ANALYSIS
+User Study.
+Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment.
+Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality.
+We collected 25 responses from an online questionnaire distributed through social media platforms.
+The questionnaire consisted of 20 groups of images along with corresponding prompts.
+Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint.
+Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest).
+As shown in Tab.
+3, our method consistently outperforms the baseline approaches.
+Specifically, by comparing the differences across various evaluation dimensions, it is evident that users showed a clear preference for our proposed method.
+This indicates that our approach achieves texture transfer results that better align with human visual perception.
+Runtime Cost.
+Runtime efficiency is crucial for 3D assets.
+Therefore, we evaluate the runtime cost of different methods on the LLFF dataset.
+As shown in Tab.
+4, our method demonstrates superior efficiency in terms of both memory usage and FPS.
+StyleGaussian incurs higher memory costs due to the use of additional VGG feature parameters during rendering.
+ARF, Ref-NPR, and SNeRF are NeRF-based methods, which result in slower rendering speeds.
+Impact of Content Loss.
+As shown in Fig.
+7, the content loss controls the strength of texture transfer, helping preserve the original scene content.
+However, an excessively large $\lambda_{c}$ hinders the scene from effectively learning the desired texture appearance.
+MORE IMPLEMENTATION DETAILS
+The frequency map is derived from high-frequency energy statistics based on the Discrete Cosine Transform (DCT).
+Specifically, according to the scale difference between the image and the feature map, the image is divided into 8×8 blocks, and DCT is performed on each block.
+The frequency density of a given region is defined as the sum of the absolute DCT coefficients within the corresponding block.
+For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings.
+In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively.
+During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}.
+For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005.
+To simplify the implementation of the geometry preservation branch, our code adopts an alternating optimization strategy between the two branches.
+MORE QUALITATIVE EVALUATION
 Texture Orientation Control.
@@
 As shown in Fig.
-6, by combining the reference texture image and view direction control, we achieve high-quality texture transfer results with controllable texture orientation.
-Table 1. Quantitative Experiment on Multi-view Consistency.
-Methods Short-term Consistency Long-term Consistency
-LPIPS($\downarrow$) RMSE($\downarrow$) LPIPS($\downarrow$) RMSE($\downarrow$)
-StyleGaussian 0.052 0.051 0.137 0.125
-ARF 0.078 0.067 0.175 0.153
-ABC-GS 0.051 0.045 0.127 0.118
-Ours 0.052 0.045 0.128 0.118
-Table 2. Quantitative Experiment on Content Protection.
-Ours ABC-GS ARF StyleGaussian
-SSIM($\uparrow$) 0.47 0.55 0.34 0.40
-CLIP-score($\uparrow$) 0.45 0.39 0.38 0.35
-5.3. Quantitative Evaluation
-In the 3D scene appearance editing task, maintaining multi-view consistency and preserving scene content information are crucial.
-To this end, we conducted extensive quantitative experiments from both aspects.
-We randomly selected 50 pairs of scene-reference image pairs from the scene and reference image datasets to quantitatively evaluate our proposed method.
-Multi-view Consistency.
-We evaluate multi-view consistency using both short-term and long-term consistency metrics.
-These metrics are computed by warping the rendered image from one viewpoint to another using optical flow (Teed and Deng, 2020).
-The RMSE and LPIPS scores are then calculated between the warped image and the target view.
-The results are shown in Table.
-1.
-It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency.
-Notably, ABC-GS disables the Gaussian densification strategy during optimization.
-In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPS.
-Content Preservation.
-For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance.
-We use SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021) to assess the preservation of content information.
-Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level.
-In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images.
-As shown in Table.
-4, our proposed method achieves significantly higher scores on both evaluation metrics.
-In particular, our method outperforms previous approaches in terms of the CLIP-score.
-It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
-Table 3. The results of user study.
-Ours ABC-GS ARF StyleGaussian
-Texture Alignment($\uparrow$) 4.23 3.02 3.08 2.15
-Visual Quality($\uparrow$) 3.90 3.37 3.18 2.22
-Figure 7. Ablation study for GT Loss.
-Figure 8. Ablation study for GTA and GPS.
-5.4. User Study
-Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment.
-Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality.
-We collected X responses from an online questionnaire distributed through social media platforms.
-The questionnaire consisted of 20 groups of images along with corresponding prompts.
-Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint.
-Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest).
-As shown in Table.
-3, our method consistently outperforms the baseline approaches.
-Specifically, by comparing the differences across various evaluation dimension, it is evident that users showed a clear preference for our proposed method.
-This indicates that our approach achieves texture transfer results that better align with human visual perception.
-5.5. Ablation Study
-Impact of GT Loss.
-GT Loss is proposed to control the texture orientation.
-In the ablation study, we perform central rotation on the same texture image to generate multiple inputs with different rotation angles.
-As shown in Fig.
-7, removing the GT Loss leads to identical texture transfer results across different input angles.
-The reason for this phenomenon is that texture images with different rotation angles produce identical texture feature sets after being processed by the GTA module.
-Without the constraint of GT Loss, the scene directly learns appearance from the aggregated texture feature set, ignoring orientation variations.
-Impact of GTA.
-As shown in Fig.
-8, removing the GTA module results in noticeable texture discontinuities and blurring in the texture transfer outputs.
-This degradation is attributed to the absence of geometry-aware texture features.
-Without the GTA module, a large portion of texture features aligned with the scene’s geometry are eliminated.
-The remaining texture features are insufficient to support learning across the full range of geometric structures within the scene.
-Learning texture features of the same orientation across different viewpoints fails to achieve continuous texture transfer results between views.
-Impact of GPS.
-The 3D texture transfer tasks inherently lack ground truth supervision, which can introduce incorrect geometry during the scene appearance editing process.
-This issue is particularly pronounced in the 3DGS-based methods, where appearance and geometry are tightly entangled.
-To address this, we propose GPS, which leverages the original content images to optimize the geometric parameters of Gaussians and correct inaccurate geometry.
-As illustrated in Fig.
-8, removing GPS results in noticeable artifacts across the scene, especially in challenging 360-degree outdoor environments.
-This highlights the critical role of GPS in maintaining geometric consistency.
-Through GPS, our texture transfer framework achieves a balance between learning texture features and preserving scene geometry.
-6. Conclusion
-In this paper, we introduced GT2-GS, a novel Geometry-aware Texture Transfer framework for Gaussian Splatting that achieves high-quality texture transfer results.
-Unlike previous 3d stylization methods, our approach explicitly considered the intrinsic relationship between geometry and texture.
-We observed that feature vectors directly extracted from texture images are insufficient to capture the rich geometric information present in 3D scenes.
-To address this, we proposed the GTA module, which enhanced the texture feature set by incorporating geometric cues such as perspective geometry, edges, and shapes.
-Furthermore, we designed a GT Loss that leveraged 3D geometric information across viewpoints to enable controllable texture transfer.
-To preserve accurate geometry during texture learning, we introduced the GPS strategy to maintain structural integrity throughout the optimization process.
-Extensive quantitative and qualitative experiments demonstrated the effectiveness of our proposed method.
-Moreover, our framework was capable of generating high-quality stylization results, showcasing its generalizability.
-Acknowledgements.
-This work is supported by the National Natural Science Foundation of China (62472178, 62376244), and Shanghai Urban Digital Transformation Special Fund Project (202301027).
-We also thank the financial support from Shanghai China Fortune Co., Ltd.
-Appendix A ADDITIONAL ANALYSIS
-Quantitative Ablation Study on GTA and GPS.
-We perform a quantitative ablation study on the Geometry-aware Texture Augmentation (GTA) module and the Geometry Preservation Strategy (GPS).
-The impact of texture transfer on the original scene content is assessed using SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021).
-We randomly selected 30 pairs of scenes and texture images for the experiment.
-As shown in Table  4, removing either of these components adversely affects the geometric structure of the scene.
-Consequently, the scene content becomes less distinguishable, leading to a decrease in semantic similarity with the original scene.
-Component-level Impact Analysis of GTA.
-We conduct a component level ablation study on the proposed GTA module, which includes multi-scale texture enhancement to simulate perspective geometry and multi-angle enhancement to match geometric cues such as edges and contours in the rendered images.
-As shown in Fig.
-9, the multi-angle enhancement significantly impacts the final texture transfer results, while removing the multi-scale enhancement leads to a lack of depth perception in the scene.
-Component-level Impact Analysis of GPS.
-GPS consists of two key components: dynamic prior injection and the geometry correction stage.
-We conduct ablation studies to investigate their impact on scene geometry.
-As shown in Fig.
-10, removing dynamic prior injection disrupts the geometric structure of textures that are difficult to learn in the specified orientation, resulting in blurry renderings.
-On the other hand, removing the geometry refinement stage leads to severe Gaussian floating artifacts throughout the scene, severely degrading the 3D geometric integrity.
-Impact of Content Loss.
-As shown in Fig.
-11, the content loss controls the strength of texture transfer, helping preserve the original scene content.
-However, an excessively large $\lambda_{c}$ hinders the scene from effectively learning the desired texture appearance.
-Impact of Color Transfer.
-We follow ARF (Zhang et al., 2022) to perform view-consistent color transfer.
-As shown in Fig.
-12, without color transfer, the scene struggles to learn colors consistent with the texture image, which significantly degrades the visual quality of 3D texture transfer.
-Appendix B MORE IMPLEMENTATION DETAILS
-For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings.
-In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively.
-During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1} and perform $N_{iter}=10$ rounds of GPS optimization, where each round consists of 80 texture transfer steps followed by 400 geometry correction steps.
-For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005 and use $N_{iter}=6$ rounds with 400 texture transfer steps and 1000 geometry correction steps per round to accommodate more viewpoints.
-Figure 9. Component-level Ablation Study of GTA.
-Table 4. Quantitative Experiment on Content Protection.
-Ours w/o GTA w/o GPS
-SSIM($\uparrow$) 0.43 0.37 0.39
-CLIP-score($\uparrow$) 0.36 0.35 0.34
-Figure 10. Component-level Ablation Study of GPS.
-Figure 11. Impact of Content Loss.
-Figure 12. Impact of Color Transfer.
-Figure 13. Qualitative Comparison of Texture Transfer on the LLFF Dataset.
-Figure 14. Qualitative Comparison of Style Transfer on the LLFF Dataset.
-Figure 15. Qualitative Comparison of Texture Transfer on the T&T Dataset.
-Figure 16. Qualitative Comparison of Style Transfer on the T&T Dataset.
-Appendix C MORE QUALITATIVE EVALUATION
-For more qualitative evaluation, we compare our method with ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024).
+8, by combining the reference texture image and view direction control, we achieve high-quality texture transfer results with controllable texture orientation.
+More Comparison.
+For more qualitative evaluation, we compare our method with SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref) and SNeRF (nguyen2022snerf).
 Fig.
-13, Fig.
-14, Fig.
-15, and Fig.
+9 and Fig.
 10 present the qualitative results.
-Appendix D DISCUSSION, LIMITATION, AND FUTURE WORK
-Our method primarily consists of three key components: Geometry-aware Texture Augmentation (GTA), Geometric-consistent Texture (GT) Loss, and Geometry Preservation Strategy (GPS).
-Among them, the proposed GTA module and GT Loss leverage 2D and 3D geometric information respectively, and their computational process is independent of specific 3D representations.
-Therefore, they can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel.
-Meanwhile, GPS provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision.
+Ours SGSST ABC-GS ARF
+Texture Alignment($\uparrow$) 4.24 3.16 3.04 3.04
+Visual Quality($\uparrow$) 3.92 3.20 3.36 3.08
+Table 3: The results of user study.
+Ours SGSST ABC-GS StyleGaussian ARF Ref-NPR SNeRF
+Memory(GB)($\downarrow$) 1.1 1.7 1.2 6.1 1.4 1.4 1.4
+FPS($\uparrow$) 151 132 133 143 8 7 7
+Table 4: Runtime Cost Comparison on the LLFF Dataset.
+Figure 7: Ablation Study of Content Loss.
+Figure 8: Qualitative Results of Texture Orientation Control.
+Figure 9: Comparison of Qualitative Results.
+Figure 10: Comparison of Qualitative Results.
+DISCUSSION AND LIMITATION
+Our method primarily consists of three key components: Geometry-aware Texture Transfer Loss (GT2 Loss), Adaptive Fine-grained Control Module (AFCM), and Geometry Preservation Branch (GPB).
+Among them, the proposed GT2 Loss and AFCM can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel.
+Meanwhile, GPB provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision.
 However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal.
-Due to the introduction of a geometry correction stage, our method introduces additional optimization time.
-Nevertheless, for forward-facing scenes, our method is able to achieve high-quality texture transfer results within 5 minutes.
+FUTURE WORK
 In future work, we plan to focus on two main aspects.
-First, we aim to enhance the geometry correction stage of GPS by integrating more accurate 3DGS optimization techniques for geometry reconstruction.
+First, we aim to enhance the geometry correction stage of GPB by integrating more accurate 3DGS optimization techniques for geometry reconstruction.
 This improves the quality of scene appearance editing.

### v4

@@
 Introduction
-3D style transfer (zhang2022arf; zhang2024stylizedgs) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information.
+3D style transfer (Zhang et al.
+2022, 2024) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information.
 With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has increased significantly.
 3D stylization techniques offer a promising solution by accelerating the creation of 3D content, particularly in complex 3D scene environments.
-Recently, the emergence of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl3Dgaussians) has significantly advanced the field of 3D stylization.
-NeRF-based stylization methods (zhang2022arf; zhang2023ref; nguyen2022snerf), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry.
-In contrast, 3DGS-based approaches (galerne2025sgsst; liu2025abc; liu2024stylegaussian) offer benefits such as explicit editability and real-time rendering.
+Recently, the emergence of Neural Radiance Fields (NeRF) (Mildenhall et al.
+2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al.
+2023) has significantly advanced the field of 3D stylization.
+NeRF-based stylization methods (Zhang et al.
+2022, 2023; Nguyen-Phuoc, Liu, and Xiao 2022), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry.
+In contrast, 3DGS-based approaches (Galerne et al.
+2025; Liu et al.
+2025, 2024) offer benefits such as explicit editability and real-time rendering.
 3D stylization methods can be broadly categorized into feed-forward and optimization-based approaches, depending on whether they support zero-shot style transfer.
-While feed-forward methods (liu2023stylerf; huang2022stylizednerf; liu2024stylegaussian) enable zero-shot stylization, they often exhibit lower rendering quality.
-In contrast, optimization-based methods (zhang2022arf; liu2025abc; zhang2024stylizedgs; galerne2025sgsst) typically produce higher-fidelity stylization results.
+While feed-forward methods (Liu et al.
+2023; Huang et al.
+2022; Liu et al.
+2024) enable zero-shot stylization, they often exhibit lower rendering quality.
+In contrast, optimization-based methods (Zhang et al.
+2022; Liu et al.
+2025; Zhang et al.
+2024; Galerne et al.
+2025) typically produce higher-fidelity stylization results.
 Despite these differences, both types of methods commonly define optimization objectives within the VGG feature space during training.
@@
 Texture Transfer
-Early texture transfer algorithms (ashikhmin2003fast; lee2010directional; hertzmann2023image; efros2023image) were based on non-parametric texture synthesis.
+Early texture transfer algorithms (Ashikhmin 2003; Lee et al.
+2010; Hertzmann et al.
+2023; Efros and Freeman 2023) were based on non-parametric texture synthesis.
 These methods preserve the structural information of the target image to achieve texture transfer effects.
 Gatys et al.
-(gatys2016image) pioneered the use of convolutional neural networks (CNNs) (simonyan2014very) for image style transfer, achieving impressive results.
-Subsequently, a variety of deep learning-based style transfer methods (chen2016fast; li2016combining; huang2017arbitrary; huo2021manifold) have been proposed.
+(Gatys, Ecker, and Bethge 2016) pioneered the use of convolutional neural networks (CNNs) (Simonyan and Zisserman 2014) for image style transfer, achieving impressive results.
+Subsequently, a variety of deep learning-based style transfer methods (Chen and Schmidt 2016; Li and Wand 2016; Huang and Belongie 2017; Huo et al.
+2021) have been proposed.
 These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer.
-Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (vaswani2017attention), some deep learning-based texture transfer methods (wang2022texture; chen2022auv; pu2024dynamic) can transfer local textures from the reference image to the target image based on high-level semantic correspondence.
+Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (Vaswani et al.
+2017), some deep learning-based texture transfer methods (Wang et al.
+2022; Chen, Yin, and Fidler 2022; Pu et al.
+2024) can transfer local textures from the reference image to the target image based on high-level semantic correspondence.
 In this work, we explore how to directly transfer texture features into 3D representations.
@@
 3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency.
-Early research in 3D scene style transfer focused on explicit representations such as point clouds (cao2020psnet; huang2021learning), meshes (hollein2022stylemesh; yin20213dstylenet), and voxels (guo2021volumetric; klehm2014property).
+Early research in 3D scene style transfer focused on explicit representations such as point clouds (Cao et al.
+2020; Huang et al.
+2021), meshes (Höllein, Johnson, and Nießner 2022; Yin et al.
+2021), and voxels (Guo et al.
+2021; Klehm et al.
+2014).
 Zhang et al.
-(zhang2022arf) explored the potential of NeRF in 3D stylization tasks.
+(Zhang et al.
+2022) explored the potential of NeRF in 3D stylization tasks.
 By optimizing the scene representation with the proposed 3D style transfer loss, it achieved pleasing visual effects.
-Subsequently, a variety of NeRF-based 3D style transfer methods (huang2022stylizednerf; pang2023locally; liu2023stylerf; zhang2024coarf; jung2024geometry) have emerged.
-Recently, the emergence of 3DGS (kerbl3Dgaussians) has brought new possibilities for scene stylization.
+Subsequently, a variety of NeRF-based 3D style transfer methods (Huang et al.
+2022; Pang, Hua, and Yeung 2023; Liu et al.
+2023; Zhang, Fernandez-Labrador, and Schroers 2024; Jung et al.
+2024) have emerged.
+Recently, the emergence of 3DGS (Kerbl et al.
+2023) has brought new possibilities for scene stylization.
 It features fast training speed, high rendering quality, and efficient performance.
-Some works (liu2024stylegaussian; zhang2024stylizedgs; mei2024regs; liu2025abc) have already explored the potential of stylization under this representation.
-Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (liu2024stylegaussian) and optimization-based approaches (zhang2024stylizedgs; mei2024regs; liu2025abc).
+Some works (Liu et al.
+2024; Zhang et al.
+2024; Mei, Xu, and Patel 2024; Liu et al.
+2025) have already explored the potential of stylization under this representation.
+Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (Liu et al.
+2024) and optimization-based approaches (Zhang et al.
+2024; Mei, Xu, and Patel 2024; Liu et al.
+2025).
 Feed-forward methods enable zero-shot style transfer.
-For instance, StyleGaussian (liu2024stylegaussian) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (huang2017arbitrary) to achieve real-time stylization.
+For instance, StyleGaussian (Liu et al.
+2024) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (Huang and Belongie 2017) to achieve real-time stylization.
 On the other hand, optimization-based methods attain higher-quality style transfer results.
-Some works (zhang2024stylizedgs; liu2025abc) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
+Some works (Zhang et al.
+2024; Liu et al.
+2025) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
 However, existing methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images.
@@
 3D Gaussian Splatting
-3D Gaussian Splatting (kerbl3Dgaussians) is a novel explicit 3D representation method.
+3D Gaussian Splatting (Kerbl et al.
+2023) is a novel explicit 3D representation method.
 It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes.
@@
 The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$.
-The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (lassner2021pulsar).
+The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (Lassner and Zollhofer 2021).
 Specifically, the Gaussians are grouped into tiles and sorted by depth.
@@
 By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized.
-Existing optimization-based 3D style transfer methods (zhang2022arf; pang2023locally; zhang2024coarf; zhang2024stylizedgs) typically adopt the nearest neighbor feature matching (NNFM) loss (zhang2022arf) as their style loss function.
+Existing optimization-based 3D style transfer methods (Zhang et al.
+2022; Pang, Hua, and Yeung 2023; Zhang, Fernandez-Labrador, and Schroers 2024; Zhang et al.
+2024) typically adopt the nearest neighbor feature matching (NNFM) loss (Zhang et al.
+2022) as their style loss function.
 It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them.
 Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$.
-The same feature extractor (e.g., VGG (simonyan2014very)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively.
+The same feature extractor (e.g., VGG (Simonyan and Zisserman 2014)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively.
 Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
@@
 Datasets.
-For the scene datasets, we utilize the LLFF dataset (mildenhall2019local) and the Tanks and Temples (T&T) dataset (knapitsch2017tanks), which are collected from the real world.
-Additionally, we use images from the ARF (zhang2022arf) style dataset and the DTD dataset (cimpoi2014describing) as reference image datasets.
+For the scene datasets, we utilize the LLFF dataset (Mildenhall et al.
+2019) and the Tanks and Temples (T&T) dataset (Knapitsch et al.
+2017), which are collected from the real world.
+Additionally, we use images from the ARF (Zhang et al.
+2022) style dataset and the DTD dataset (Cimpoi et al.
+2014) as reference image datasets.
 Baseline.
-We compare our method with the state-of-the-art 3D stylization methods, including SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref), and SNeRF (nguyen2022snerf).
+We compare our method with the state-of-the-art 3D stylization methods, including SGSST (Galerne et al.
+2025), ABC-GS (Liu et al.
+2025), StyleGaussian (Liu et al.
+2024), ARF (Zhang et al.
+2022), Ref-NPR (Zhang et al.
+2023), and SNeRF (Nguyen-Phuoc, Liu, and Xiao 2022).
 Specifically, SGSST, ABC-GS, and StyleGaussian are based on 3DGS, and ARF, Ref-NPR, and SNeRF are based on NeRF.
@@
 Implementation Details
-We perform view-consistent color transfer (zhang2022arf) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters.
-For the VGG (simonyan2014very) feature extractor, we employ the conv3 block of VGG-16.
+We perform view-consistent color transfer (Zhang et al.
+2022) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters.
+For the VGG (Simonyan and Zisserman 2014) feature extractor, we employ the conv3 block of VGG-16.
 When associating color parameters with texture features, we set the default number of depth groups $K$ to 4, and the rotation angle $\theta$ is sampled across the full 360 degrees.
@@
 Multi-view Consistency.
-We evaluate multi-view consistency (liu2024stylegaussian) using both short-term and long-term consistency metrics.
+We evaluate multi-view consistency (Liu et al.
+2024) using both short-term and long-term consistency metrics.
 The results are shown in Tab.
@@
 For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance.
-We use SSIM (wang2004image) and CLIP-score (radford2021learning) to assess the preservation of content information.
+We use SSIM (Wang et al.
+2004) and CLIP-score (Radford et al.
+2021) to assess the preservation of content information.
 Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level.
@@
 Meanwhile, we thank the financial support from Shanghai Chinafortune Co., Ltd.
-ADDITIONAL ANALYSIS
-User Study.
-Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment.
-Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality.
-We collected 25 responses from an online questionnaire distributed through social media platforms.
-The questionnaire consisted of 20 groups of images along with corresponding prompts.
-Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint.
-Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest).
-As shown in Tab.
-3, our method consistently outperforms the baseline approaches.
-Specifically, by comparing the differences across various evaluation dimensions, it is evident that users showed a clear preference for our proposed method.
-This indicates that our approach achieves texture transfer results that better align with human visual perception.
-Runtime Cost.
-Runtime efficiency is crucial for 3D assets.
-Therefore, we evaluate the runtime cost of different methods on the LLFF dataset.
-As shown in Tab.
-4, our method demonstrates superior efficiency in terms of both memory usage and FPS.
-StyleGaussian incurs higher memory costs due to the use of additional VGG feature parameters during rendering.
-ARF, Ref-NPR, and SNeRF are NeRF-based methods, which result in slower rendering speeds.
-Impact of Content Loss.
-As shown in Fig.
-7, the content loss controls the strength of texture transfer, helping preserve the original scene content.
-However, an excessively large $\lambda_{c}$ hinders the scene from effectively learning the desired texture appearance.
-MORE IMPLEMENTATION DETAILS
-The frequency map is derived from high-frequency energy statistics based on the Discrete Cosine Transform (DCT).
-Specifically, according to the scale difference between the image and the feature map, the image is divided into 8×8 blocks, and DCT is performed on each block.
-The frequency density of a given region is defined as the sum of the absolute DCT coefficients within the corresponding block.
-For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings.
-In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively.
-During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}.
-For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005.
-To simplify the implementation of the geometry preservation branch, our code adopts an alternating optimization strategy between the two branches.
-MORE QUALITATIVE EVALUATION
-Texture Orientation Control.
-The proposed GT Loss enables the scene to propagate the target texture information based on the prior view.
-Therefore, by controlling the learnable texture direction in the initial prior view, we can influence the overall appearance learned by the entire scene.
-This control is achieved by injecting pseudo prior angle information into the first prior view.
-As shown in Fig.
-8, by combining the reference texture image and view direction control, we achieve high-quality texture transfer results with controllable texture orientation.
-More Comparison.
-For more qualitative evaluation, we compare our method with SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref) and SNeRF (nguyen2022snerf).
-Fig.
-9 and Fig.
-10 present the qualitative results.
-Ours SGSST ABC-GS ARF
-Texture Alignment($\uparrow$) 4.24 3.16 3.04 3.04
-Visual Quality($\uparrow$) 3.92 3.20 3.36 3.08
-Table 3: The results of user study.
-Ours SGSST ABC-GS StyleGaussian ARF Ref-NPR SNeRF
-Memory(GB)($\downarrow$) 1.1 1.7 1.2 6.1 1.4 1.4 1.4
-FPS($\uparrow$) 151 132 133 143 8 7 7
-Table 4: Runtime Cost Comparison on the LLFF Dataset.
-Figure 7: Ablation Study of Content Loss.
-Figure 8: Qualitative Results of Texture Orientation Control.
-Figure 9: Comparison of Qualitative Results.
-Figure 10: Comparison of Qualitative Results.
-DISCUSSION AND LIMITATION
-Our method primarily consists of three key components: Geometry-aware Texture Transfer Loss (GT2 Loss), Adaptive Fine-grained Control Module (AFCM), and Geometry Preservation Branch (GPB).
-Among them, the proposed GT2 Loss and AFCM can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel.
-Meanwhile, GPB provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision.
-However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal.
-FUTURE WORK
-In future work, we plan to focus on two main aspects.
-First, we aim to enhance the geometry correction stage of GPB by integrating more accurate 3DGS optimization techniques for geometry reconstruction.
-This improves the quality of scene appearance editing.
-In addition, we will explore the integration of diffusion models and multimodal large language models (MLLM) with 3D texture and style transfer.
-Leveraging the rich prior knowledge of multimodal models, we envision a unified framework that supports both text-driven and image-driven style transfer in 3D scenes.